### v1

# Exposía: Academic Writing Assessment of Exposés and Peer Feedback

Abstract
We present Exposía, the first public dataset [^1] https://tudatalib.ulb.tu-darmstadt.de/handle/tudatalib/4978 that connects writing and feedback assessment in higher education, enabling research on educationally grounded approaches to academic writing evaluation. Exposía includes student research project proposals and peer and instructor feedback consisting of comments and free-text reviews. The dataset was collected in the “Introduction to Scientific Work” course of the Computer Science undergraduate program that focuses on teaching academic writing skills and providing peer feedback on academic writing. Exposía reflects the multi-stage nature of the academic writing process that includes drafting, providing and receiving feedback, and revising the writing based on the feedback received. Both the project proposals and peer feedback are accompanied by human assessment scores based on a fine-grained, pedagogically-grounded schema for writing and feedback assessment that we develop.
We use Exposía to benchmark state-of-the-art open-source large language models (LLMs) for two tasks: automated scoring of (1) the proposals and (2) the student reviews. The strongest LLMs attain high agreement on scoring aspects that require little domain knowledge but degrade on dimensions evaluating content, in line with human agreement values. We find that LLMs align better with the human instructors giving high scores. Finally, we establish that a prompting strategy that scores multiple aspects of the writing together is the most effective, an important finding for classroom deployment.
Exposía: Academic Writing Assessment of Exposés and Peer Feedback
Dennis Zyska1, Alla Rozovskaya2, Ilia Kuznetsov1, Iryna Gurevych1 1Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science and Hessian Center for AI (hessian.AI), Technical University of Darmstadt 2Department of Computer Science at Queens College, City University of New York (CUNY) 1www.ukp.tu-darmstadt.de
1 Introduction
Figure 1: Overview of Exposía. In the university course “Introduction to Scientific Work”, students submit a draft exposé, receive feedback in the form of comments and a free-text review, revise and submit a final exposé. Feedback is produced by peers and instructors. Instructors grade draft and final exposés and student feedback.
Writing a research project proposal [^2] We will use research project proposal, research project exposé, or exposé interchangeably in the paper. is the number one challenge in academic education and represents a pressing concern across the academic community Muneer et al. (2020). It is also an essential skill required for conducting research and for achieving academic success at all levels Kivunja (2016); Grech (2018). However, a lot of students lack the necessary competencies to develop a research proposal and its core components, including identifying a research question, formulating appropriate research methodology, and defining hypotheses and methods to solve the problem Rastri et al. (2023).
At the same time, higher education and research in general face another pressing challenge: the peer review crisis Lu et al. (2025). High-quality feedback is an essential pedagogical tool that has been shown to vastly improve learning outcomes of both authors and reviewers Wu and Schunn (2021). Yet, students are rarely taught this skill in a classroom setting. Together, these challenges highlight the need for instructional approaches that simultaneously support research project proposal writing and developing effective peer review skills.
Teaching students how to write an effective project exposé and how to provide high-quality feedback is crucial in university education for several reasons. First, assessing both project exposés and peer feedback requires substantial human effort and does not scale well Grabau and Wilson (1995). Second, developing automated assessment tools requires modeling complex cross-document relationships, connecting student writing, feedback, and revisions. Moreover, peer feedback is often of low quality and highly inconsistent Purkayastha et al. (2025). Finally, collecting data in an educational setting is extremely challenging (Section 3). As a result, existing resources for computational writing assessment are fragmented across subtasks such as automated essay scoring Ke and Ng (2019), peer feedback analysis Staudinger et al. (2024), or revision modeling Ruan et al. (2024).
In this work, we present Exposía, the first corpus that integrates drafting, feedback, revision, and fine-grained exposé and feedback assessment scores within a single dataset (Figure 1). Exposía is continuously collected in a newly developed university Computer Science course. The course focuses on teaching students (1) how to write a research project exposé and (2) how to write peer feedback. While this is an undergraduate course, the skills it develops are equally applicable to graduate students and other early-career researchers.
Exposía captures a complete research-proposal writing workflow. As illustrated in Figure 1, each student first produces an exposé draft. The draft receives (i) peer feedback from fellow students and (ii) instructor feedback, both provided as (a) inline comments anchored to draft passages and (b) a free-text review. The student revises the draft based on the feedback to produce the final exposé version. Figure 2 (top) details each component.
Figure 2: Example instance from Exposía. Top: A student draft exposé (top-left) is reviewed by an instructor or a student peer and receives feedback. The feedback consists of (1) passage-anchored inline comments (shown in the top middle in red and green), each associated with a tag (e.g., Weakness, Strength); and (2) a free-text review (top-right). Bottom: Exposés and reviews are graded by instructors and receive scores on multiple aspects (criteria) organized into high-level topics (rubrics). The exposé grading criteria and the review grading criteria are shown in the bottom left and bottom right of the figure, respectively.
We further design fine-grained grading criteria that target the key aspects of exposé and feedback writing (Figure 2, bottom). These criteria are grounded in established research on exposé writing and feedback assessment and are tailored to the academic disciplinary context of Exposía. The exposé drafts, the final exposés, and the student feedback are graded by course instructors, who provide scores on each of the criteria.
To the best of our knowledge, Exposía is the first dataset, collected in an educational setting, that features this unique class of cross-document relations, enabling new applications for writing and grading assistance. Because the corpus contains paired drafts and final exposés, it enables quantifying draft-to-final improvements on specific criteria and relating these changes to the feedback received. Overall, we find that exposés improve in response to feedback, particularly on criteria related to the in-domain content of the writing rather than form. We further establish baseline models with open-source LLMs for two tasks: exposé scoring and feedback scoring, based on the fine-grained grading schema for writing and feedback assessment.
Contributions
The main contribution of this work is the introduction of a novel problem definition of exposé and review scoring in the educational context. The result is the first public dataset of student research exposés and feedback, accompanied by human assessment scores on exposés and feedback, based on the fine-grained grading schema that we develop. We present in-depth analyses of the dataset, and models for the corresponding tasks of exposé and feedback scoring, establishing baselines that the community can build on.
2 Related Work
Automated Essay Scoring
The task of Automated Essay Scoring (AES) Page (1966), aims to reduce human grading effort and improve consistency in writing assessment. Most widely used AES benchmarks consist of essays written to standardized prompts in general-domain language. These essays are typically manually annotated with a holistic score, complemented by scores for a small set of assessment criteria (e.g., content, organization, language) Mathias and Bhattacharyya (2018); Gaudeau (2025); Li and Ng (2024); Xue et al. (2021); Crossley et al. (2023); Ishikawa (2023). Importantly, these benchmarks do not include human feedback, revisions, or feedback assessment scores.
[…]
LLM-based scoring and transfer to peer review Recent work evaluates LLMs as scorers and emphasizes that educational use requires evidence beyond accuracy Pack et al. (2024); Huang and Wilson (2025). Exposía complements this line of work, by enabling fine-grained exposé and feedback assessment with LLMs and interpreting LLM–human alignment relative to human agreement (Section 6). Because our corpus links each exposé draft with feedback text and assessment scores, it is related to scholarly peer-review datasets Kang et al. (2018); Dycke et al. (2023). However, these datasets provide only coarse-grained annotations. Importantly, they serve fundamentally different purposes, as they focus on scientific communication in a professional decision-making context, whereas Exposía is situated in an educational environment with a focus on learning and assessment. This creates a testbed for transfer experiments between educational feedback assessment and scholarly reviewing.
3 Course Setting and Study Protocol
Data collection protocol
The dataset was collected in the newly-introduced Bachelor-level Computer Science course Introduction to Scientific Work offered for the first time in winter term 2024. The corpus is the result of a targeted data collection. [^3] The institutional IRB has approved the related proposal. Collecting data in an educational setting is challenging due to the high sensitivity of the data: collection, storage, and sharing of educational data raise serious privacy and ethical issues Guan et al. (2023).
Course structure The students wrote an exposé on one of three topics, covering various areas of Computer Science and Artificial Intelligence (AI): [^4] Topics were evenly assigned based on preference voting. Holographic AI Assistant (Human-Computer Interaction), Interactive LLM-based Teaching Agents (Natural Language Processing), or Generative AI in Healthcare (Data Mining).
In the beginning of the course, [^5] The syllabus is provided in Appendix A. alongside the lectures, the students received guidelines for writing an exposé (Appendix B) and the exposé grading criteria (Section 4.2). Each student produced an exposé draft. This was followed by a peer-review phase, where each student wrote double-blind reviews for two exposés written by their peers, selected uniformly at random. The students were provided with review grading criteria (Section 4.3). The instructional staff also provided a review for each exposé. This was followed by a revision phase to produce the final exposé, based on the feedback. Both the drafts and final versions, and the student reviews were graded by the instructional staff. The final grade was calculated as follows: draft exposé 30%, peer reviews 30%, and final exposé 40%. [^6] Grades for each component are derived from criterion-based scores (see Section 4 and Appendix F.3).
Reviewing was facilitated through the open-access CARE platform Zyska et al. (2023, 2025).
Continued large-scale data collection 193 students enrolled in the course, [^7] Enrollment was limited because the course was offered for the first time as a pilot. of whom 159 completed all requirements. The Exposía dataset refers to the open data that participants agreed to contribute for public release. All 12 instructors and about 30% of students consented to share their materials.
Exposía currently comprises 55 exposés, 306 reviews, and 2,253 inline comments.
We emphasize that the dataset will continue to grow, as the course will be offered annually. We will continue to expand and update the dataset with additional data. In the winter term 2025, 600 students enrolled in the course. We expect at least 30% of the students to consent to the release of their materials, consistent with the previous semester.
In general, 600 students are expected to enroll each year, with a higher proportion of students willing to contribute their content, as they understand the significance of data collection in the course that teaches important skills for writing scientific papers and for providing effective feedback on scientific writing. This is consistent with similar important collection efforts (Dycke et al. (2023)).
4 The Exposía Dataset
4.1 Exposía Dataset Overview
Figure 2 illustrates two main components of the Exposía: exposés and feedback.
Exposés Each exposé is available in two versions: a draft and a revised final version produced after receiving feedback. The exposés follow a standard Computer Science research proposal template with the following sections: (1) Motivation; (2) Approach divided into three parts – Related Work, Theoretical Framework, and Methodology; (3) a Schedule outlining the week-by-week plan; and (4) a Bibliography (Appendix B).
Feedback: inline comments and reviews Each draft exposé received feedback from instructors and other students. Each feedback consists of inline comments and a free-text review. An inline comment consists of a span-level highlight on the draft text and a free-text comment connected to it (Figure 2). Each inline comment is also associated with a tag from a fixed taxonomy $\mathcal{Y}=\{\textit{Strength},\textit{Weakness},\textit{Highlight},\textit{Other}\}$. A reviewer first adds inline comments. Then the reviewer uses them as the basis for composing a review, a document-level free-text feedback on the draft. Appendix C provides more detail.
4.2 Exposé Assessment Criteria and Rubrics
Both the draft and the final exposé are graded independently by an instructional staff member on 36 criteria and include 36 per-criterion assessment scores (Appendix Table LABEL:tab:expose_rubric). Each criterion is rated on an ordinal scale of [0,1], [0,1,2], or [0,-1,-2].
Designing exposé assessment criteria and rubrics The extensive criterion-based grading scheme is designed to support exposé assessment across multiple aspects. These criteria were derived using 18 guidance documents for writing thesis exposés from major universities in Germany and Austria (Appendix E). These were supplemented with over 100 exposés previously submitted in the Computer Science Department. The 36 criteria center on the different aspects of writing a scientific proposal, with a focus on content assessment.
Guided by research in rubric construction and fairness-oriented assessment Jonsson and Svingby (2007); Mayring (2000); Kuckartz and Rädiker (2023), we organized the criteria into nine high-level categories referred to as rubrics (Table 1) that decompose evaluation into orthogonal dimensions, by grouping criteria that evaluate related aspects of the exposé writing.
Rubrics are an effective assessment tool and are constructed to ensure consistency in evaluation across a diverse student body and multiple evaluators Northern Illinois University Center for Innovative Teaching and Learning (2012). See Appendix D.1 for more details.
Code Rubric Description #Crit.
Exposé
LA Lang./Quality Assessment of language quality and structural coherence of the text. 2
ME Metadata Evaluation of title appropriateness and creativity. 2
ST Form/Structure Evaluation of document formatting and structural requirements. 2
MO Motivation Assessment of problem identification, context establishment, and research question formulation. 7
AP Approach Assessment of state of the art analysis, theoretical framework. 8
MD Methodology Assessment of the methodology. 7
SC Schedule Assessment of project timeline structure, completeness, and realism. 4
BI Bibliography Assessment of bibliography consistency and literature relevance. 2
AD Addt’l points Add points for great submissions or deduct points for major mistakes. 2
Review
SC Structure/Clarity Assessment of overall structure and clarity. 2
LG Language Assessment of language quality and tone. 2
FU Feed Up Assessment of learning goals and success criteria. 2
FB Feed Back Assessment of feedback quality and knowledge transfer. 6
$F$ Feed Forward Assessment of future learning guidance. 2
CQ Content Quality Assessment of content accuracy and usefulness. 3
ER Errors Assessment of potential errors and issues in feedback. 5
AD Addt’l points Add points for great reviews or deduct points for major mistakes. 2
Table 1: Exposé and review grading rubrics. The last column shows the number of criteria in the rubric.
4.3 Review Assessment Criteria and Rubrics
Each student review was independently graded by an instructional staff member and received 24 assessment scores, in line with the 24-criterion grading scheme (Appendix Table LABEL:tab:review_rubric) organized into eight rubrics (Table 1), with each criterion rated on an ordinal scale of [0,1], [0,1,2], or [-1,1]. The rubrics are shown in Table 1.
Designing review assessment criteria and rubrics The main goal of assessing reviews was to capture the extent to which peer feedback supports learning, reflection, and revision. Accordingly, the rubrics operationalize key educational principles of effective feedback—clarity, constructiveness, specificity, and usefulness—through an analytic, criterion-based framework that is transparent, fair, and aligned with established feedback theory Hattie and Timperley (2007). As such, our rubric design is primarily informed by the feedback model in Hattie and Timperley (2007), which conceptualizes feedback as answering three questions: Where am I going? (Feed Up), How am I going? (Feed Back), and Where to next? (Feed Forward).
Drawing on Narciss (2008), we integrated the components of feedback as criteria within the 8 rubrics. We further incorporated insights from D’Arcy et al. (2024), which analyzes revisions in response to peer reviews in scientific writing, highlighting that actionable comments, those offering concrete, suggestions, are more likely to lead to substantive text revisions. As such, our rubrics emphasize argumentation, accuracy, and clear suggestions for improvement as indicators of pedagogically effective feedback Shute (2008).
Finally, while actionable feedback promotes learning, ineffective or misleading reviews can have the opposite effect. To this end, inspired by Du et al. (2024), the criteria include a penalty for unhelpful feedback patterns (e.g., neglect of key elements, unsupported claims, or contradictory advice). Appendix D.2 provides more detail.
5 Exposía Dataset Analysis
5.1 Exposía Dataset Statistics
The dataset statistics are in Table 2. As the dataset comprises only materials for which consent was granted, some exposés lack the full set of three reviews (Table 18), and there exist 184 orphaned reviews because the corresponding exposés could not be made public. The total number of student reviews is 110, including 34 reviews with an existing exposé and 76 that are orphaned. Each draft exposé includes on average 41 inline comments. Weakness is the most frequent tag (59%) of all comments, followed by Strength (18%), Other (14%) and Highlight (9%).
Topic Exposés IC Reviews
T1 19 798 43
T2 18 785 44
T3 18 670 35
Orphaned reviews* 184
Total $\Sigma$ 55 2,253 306
Table 2: Statistics on the Exposía corpus. Topics: T1=The Holographic AI Assistant, T2=Interactive LLM-based teaching agents, T3=Generative AI in Healthcare. IC=Inline Comments. *Orphaned reviews counts reviews for which no corresponding exposé exists.
Exposé revision analysis We analyze the impact of feedback by comparing how the final exposé changes relative to the draft: final versions increase in length from 528 to 649 words, on average. We also analyze revisions with respect to exposé criterion-based scores, averaged by rubric. The largest gains occur in Approach, Motivation, Language/Quality and Methodology (Figure 3). The most substantial revisions focus on conceptual and argumentative elements (see Appendix F.1).
Figure 3: Per-rubric score improvements from draft to final exposé. Each bar shows the mean change in score. Error bars indicate nonparametric bootstrap 95% confidence intervals for each item.
Figure 4: Human–LLM agreement (QWA) for exposé scoring by expertise level. The dotted line shows human-human QWA.
Figure 5: Group asymmetry in human–LLM agreement by exposé rubric. Each point shows the difference in agreement between an LLM and human raters (Group 1 (G1) vs. Group 2 (G2)). The vertical line at $\Delta=0$ indicates equal agreement of the LLM with human G1 and G2 raters. Points to the right indicate the model agrees more with G1 raters than with G2 raters.
Review analysis We compare reviews written by instructors vs. students (Appendix F.2): instructor reviews are almost twice as short as student reviews, but have higher lexical density and higher lexical diversity. Appendix F.2 provides more analyses.
5.2 Inter-Annotator Agreement
Rubric QWA $\boldsymbol{\alpha}$ $\mathbf{r}$
Exposé rubrics
Language/Quality (#2) 0.85 0.10 0.19
Metadata (#2) 0.83 0.00 0.31
Form/Structure (#2) 0.97 0.33 0.34
Motivation (#7) 0.76 0.11 0.17
Approach (#8) 0.77 0.22 0.26
Methodology (#7) 0.62 0.18 0.22
Schedule (#4) 0.93 0.36 0.44
Bibliography (#2) 0.78 0.19 0.42
Addt’l points (#2) 0.98 0.32 0.34
All (avg.) 0.79 0.20 0.27
Review rubrics
Structure/Clarity (#2) 0.86 0.27 0.38
Language (#2) 0.93 0.09 0.17
Feed Up (#2) 0.81 0.43 0.43
Feed Back (#6) 0.86 0.04 0.08
Feed Forward (#2) 0.78 0.02 0.04
Content Quality (#3) 0.94 0.22 0.24
Errors (#5) 0.93 0.10 0.12
Addt’l points (#2) 0.98 -0.01
All (avg.) 0.89 0.13 0.18
Table 3: IAA on exposé and review scoring. Quadratic weighted agreement (QWA), Krippendorff’s ($\boldsymbol{\alpha}$), and Pearson $r$. $r$ for Addt’l points (review) is undefined due to zero variance; we exclude it from the average. Result for each rubric is averaged over all criteria in the rubric. The number of criteria per rubric is next to rubric name.
We compute inter-annotator agreement (IAA) on exposé and review criterion-based scores. Each draft exposé and student review was independently graded by two instructors on 36 exposé and 24 review scoring criteria, respectively.
Rater training During the semester, each draft exposé and student review was graded by an instructor only once; after the course, each item was graded again for calculating IAA. We refer to instructors who graded during the semester as Group 1. They attended weekly sessions and discussed edge cases. Group 1 raters were instructed to resolve uncertainty in favor of the student, because grades were at stake. Group 2 includes a different subset of instructors (with one who also participated in Group 1). They received the same guidelines, but they were not instructed to resolve uncertainty in favor of the student. See Appendix G.1 for more details.
IAA metrics We report Krippendorff’s $\alpha$ Krippendorff (2019), [^8] Krippendorff’s $\alpha$ can be used in a setup such as ours, where we are comparing multiple raters where raters also belong to one of the two groups. Pearson correlation coefficient $r$, and raw quadratic weighted percent agreement (QWA). We have several criteria with skewed distributions where the majority label accounts for over 90% cases (Tables 19 and  20). In this case, both $\alpha$ and $r$ can be misleading Jeni et al. (2013); Zhao et al. (2022), and QWA is the most informative metric. Appendix G.2 and G.3 provide more detail.
IAA on exposé assessment scores Table 3 (top) summarizes IAA results by exposé rubric. Overall agreement yields an average QWA of $0.787$, with $\alpha=0.198$, $r=0.270$. The observed correlation levels are comparable to those reported on automated essay scoring Persing and Ng (2014, 2013) and reflect the subjectivity of the task. The rubrics that attain the lowest agreement levels are those related to content (Motivation, Approach, Methodology). Table 19 reports IAA for all 36 criteria.
IAA on review assessment scores Table 3 (bottom) summarizes IAA results by review rubric, computed over student reviews. While QWA is higher for review rubrics than for exposé, $\alpha$ and Pearson coefficient are lower, due to the skewed distributions of many criteria (Appendix G.3). For criterion-level agreement and discussion, see Appendix G.2 and Appendix Table 20.
6 Experiments
6.1 Experimental Setup
Tasks
We use Exposía for benchmarking performance on two tasks: (i) criterion-based exposé assessment and (ii) criterion-based review assessment. Models predict scores for individual criteria.
LLM prompting We assess LLM performance in a zero-shot setting. We evaluate four state-of-the-art open-weight instruction-tuned LLMs with different architectures and sizes (Appendix H). For both tasks, we use a unified prompting strategy and compare two prompting variants: (i) single-criterion prompts, which present and score one criterion at a time, and (ii) combined prompts, which list all criteria for the respective task in a single call. Cross-criterion synergy could potentially benefit the model, with a lower execution cost. Appendix I provides more detail about the prompts.
Human Qwen3 (80B) Llama 3.3 (70B) GPT OSS (120B) GPT OSS (20B)
Exp. #Crit. $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$
Exposé criterion-based scoring (over 31 criteria)
0 3 0.89   0.91   0.85   0.90   0.90   0.75   0.72   0.73   0.81
1 20 0.79   0.76   0.79   0.79   0.78   0.81   0.78   0.82   0.78
2 8 0.64   0.57   0.62   0.63   0.62   0.68   0.69   0.68   0.69
All 31 0.76   0.72   0.75   0.76   0.75   0.77   0.75   0.77   0.76
Review criterion-based scoring (over 22 criteria)
1 17 0.89   0.89   0.85   0.91   0.84   0.82   0.75   0.77   0.71
2 5 0.86   0.87   0.82   0.86   0.77   0.67   0.63   0.64   0.63
All 22 0.88   0.89   0.84   0.90   0.82   0.78   0.72   0.74   0.69
Table 4: Quadratic weighted agreement (QWA) for the two tasks, grouped by expertise level (3 and 2 expertise levels for exposé and review criteria, respectively). Exp.=expertise level; #Crit.=number of criteria in the subset. Prompting variants: all (combined) and ind. (individual). Each LLM column shows human-LLM QWA averaged over both human raters and all criteria. Human column shows agreement between the human raters.
Evaluation Following studies on criterion-based scoring Pack et al. (2024); Huang and Wilson (2025), we evaluate performance, by comparing scores predicted with an LLM with a score assigned by a human rater, using QWA (Section 5.2). Results are reported on 55 exposés drafts and 110 student reviews. Unless otherwise specified, we compute QWA against Group 1 and Group 2 raters and average the results. For each exposé and review, a corresponding model predicts 31 and 22 scores, respectively,9995 exposé and 2 review criteria are excluded (Appendix H). for a total of 1705 exposé criteria scores and 2420 review criteria scores.
6.2 Key Results
Table 4 shows main results on exposé and review assessment (averaged over all criteria – rows All).
Single-criterion vs. combined prompting Across tasks and models, combined prompting yields higher agreement, while being more computationally efficient. The only exception is Qwen3 model on exposé scoring. Tables 22 and 23 tabulate criterion- and rubric-level results.
Performance by expertise level We further group criteria by expertise level: three levels for exposé, and two for review, based on the degree of domain knowledge required for assessment (Appendix D). Figure 4 demonstrates that exposé criteria requiring more in-domain knowledge display more variability, and, importantly, this aligns with human raters. Appendix Figure 7 reveals that for review scoring, the difference is not as dramatic, suggesting that review assessment does not require as much domain knowledge. See Appendix J for further discussion.
6.3 Additional Analyses
Systematic group bias
We observe a tendency for Group 1 raters to assign higher scores (Appendix G). This matters for interpreting LLM–human agreement: LLM agreement with Group 1 is often higher than agreement with Group 2, consistent with LLMs being more generous (Figure 5). We do not observe similar behavior for review scoring (Appendix Figure 8). This is consistent with the more dramatic differences on expertise level for exposé scoring and suggests that review assessment does not require as much in-domain expertise.
Scalability and classroom deployment Combined prompting is both more effective and more efficient computationally (Appendix J and K). This is important because student activity is strongly deadline-driven (Figure 6). Thus, combined prompting is recommended for classroom deployment.
7 Discussion
Interpreting model performance relative to human raters Human agreement varies substantially across criteria and expertise levels, indicating that some aspects of academic writing and feedback are inherently more difficult to assess consistently. Within this context, LLMs achieve agreement levels that approach the consensus of human agreement. Notably, performance degrades on expert-level criteria, mirroring the same dimensions where human agreement is weakest.
LLM alignment with human groups The tendency of LLMs to align more closely with Group 1 raters suggests that LLMs adopt a more generous grading regime Flodén (2024), likely to originate from the positive bias in LLMs Jain et al. (2025).
Potential tasks enabled by Exposía Beyond the tasks studied here, Exposía enables a range of research directions. These include modeling draft-feedback-revision links across documents, with numerous applications, such as feedback text generation. More broadly, the dataset supports studying how feedback mediates learning and writing development in higher-education settings.
8 Conclusion
We introduce a novel task definition for exposé and review scoring in an educational context and present the first public dataset, Exposía, of student research exposés and feedback, accompanied by human assessment scores that are based on the fine-grained grading schema we develop.
By integrating drafting, feedback, and revision into a single resource, Exposía enables pedagogically grounded research on writing and feedback in higher education and provides a unified framework that allows for a joint study of feedback and exposé writing quality. Although the dataset is collected in the Computer Science domain, the task formulation and the assessment framework are general enough to inform the development of assessment models in other areas.
We establish baseline performance with state-of-the-art LLMs: models achieve high agreement with human raters on criteria that do not require in-domain expertise, while higher expert-level criteria remain challenging. Our findings provide insight and recommendations for classroom deployment.
Limitations
This study has several limitations. First, Exposía is shaped by the course consent protocol: since only consented materials are included, not all exposés retain the complete set of three reviews. For the same reason, some of the reviews in Exposía are orphaned, as the author of the corresponding exposé did not provide consent (Section 5.1). As such, the consent structure reduces the number of complete draft–review–revision instances in the released data. We expect these constraints to diminish as the course is offered repeatedly and the dataset continues to grow (Section 3). Although the current size of Exposía is modest compared to some other NLP datasets, it is substantial for the educational domain, particularly given its complexity and the richness of its annotations. We discuss the time investment to create the various Exposía components (feedback and criterion-based scores on exposés and reviews) in Appendix F.3.
Second, although the course covers three areas – Human–Computer Interaction, Natural Language Processing, and Data Mining – all topics remain within Computer Science and AI. However, we believe that the task formulation and the assessment framework are general enough and can inform the development of assessment models in other areas.
Third, our model comparison prioritizes reproducibility and transparency in an educational setting and therefore focuses on openly available models. As a result, we do not evaluate proprietary closed-source LLMs, which may behave differently and limit the scope of our model comparison.
Finally, and most importantly, although LLM-based scoring can be attractive for reducing instructional workload, our results should not be interpreted as support for fully automated grading: the models exhibit systematic scoring tendencies and may miss substantive issues that are crucial for guiding revision, and automated grading can incentivize strategic “gaming” rather than genuine improvement Alqahtani et al. (2023); we therefore view LLMs primarily as decision-support tools, and as common in AES, recommend a human-in-the-loop workflow when using them for assessment.
Ethical Considerations
The data collection was approved by the University’s Institutional Review Board and Data Protection Office under IRB proposal EK 77/2024. Data was collected based on informed consent (all instructors and 55 students), following a three-tier structure: (1) general consent for data processing within the review platform, (2) optional consent for contributing textual and highlight data, and (3) optional consent for contributing interaction data. Participation in the study was strictly voluntary and had no effect on grading. No monetary compensation was provided to students, as participation was part of the course setting; importantly, contributing research data was optional. Instructors were compensated via regular University employment contracts, following institutional pay scales and policies. The compensation was not contingent on participation in the study.
We release Exposía to enable research on academic writing, peer feedback, and assessment in educational settings. We stress that our findings are NOT to be interpreted as justification for fully automatic grading: all assessment-oriented applications of AI should be restricted to decision-support roles with human oversight, in line with the limitations outlined above.
Acknowledgments
This work has been funded by the German Research Foundation (DFG) as part of the PEER project (grant GU 798/28-1) and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We thank Dr. Ekaterina Kochmar (Mohamed bin Zayed University of Artificial Intelligence) and Dr. Martin Greisel (University of Augsburg) for their valuable feedback and discussions.
Appendix A Course Syllabus
This appendix documents the syllabus of the undergraduate computer science course Introduction to Scientific Work, which provided the instructional setting for the data collection described in Section 3.
The course covers foundational concepts and practical skills required for conducting scientific research. Here we present the weekly lectures:
  1. 1.
    What is science? Course overview and introduction
  2. 2.
    Empirical methods and research design
  3. 3.
    Literature review and academic search strategies
  4. 4.
    Project proposals (exposés); start writing exposé draft
  5. 5.
    Scientific writing
  6. 6.
    for academic writing
  7. 7.
    Research Questions and practical exercise on academic literature search
  8. 8.
    Introduction to peer review
  9. 9.
    Feedback; start writing peer feedback
  10. 10.
    Generative AI in scientific work
  11. 11.
    Research ethics I
  12. 12.
    Research ethics II
  13. 13.
    Workshop: Software development
  14. 14.
    Workshop: Research and project management
Appendix B Exposé Template and Writing Guidelines
Panel Exposé template shows the content of the exposé template that students received as part of the course and used to create their exposés. While the complete instructions were provided in an extensive slide-based lecture (Appendix A), the excerpt reproduced here is the concise reference sheet embedded directly in the template. Because this template guided how students organized and formatted their drafts, we include it to ensure transparency and to document the pedagogical context in which the dataset was created.
Exposé template
1 Motivation
Hook – Introduce the topic and catch the reader’s interest. Concrete: Quote, example, real-world event. Helpful: Directly useful or valuable information. Convincing: Strong argument or empirical fact. Shocking: Unexpected or counterintuitive finding. Personal: Research journey or applied experience. Understandable: Link complex ideas to everyday life. Anchor – Situate the problem concisely. Discipline, field, or domain. How the problem has been addressed so far. Current academic relevance. Deficits or criticism in existing literature. Teaser – Identify the gap and position your research. Research Questions Break the main RQ into meaningful sub-questions. Clarify terminology; define if needed. Set clear boundaries and scope.
2 Approach
State of the Art (SOTA) Refer to relevant existing work. Highlight research deficits and criticism. Combine different traditions or perspectives. Theoretical Framework Introduce theories used in the study. Explain key concepts and definitions. Describe how the theory informs the RQs. Note: Theories need not be SOTA but should be connected to it. Methodology Specify method (quantitative, qualitative, mixed, source research). Define target population and sampling. State data sources and availability. Clarify agreements or permissions. Anticipate potential difficulties. Discuss methodological limitations. Explain relevance and scope. Describe variables and their relationships. Outline practical implementation (software, documents, helpers). Clarify epistemic goals (knowledge to be gained). If building on prior work, state progress so far.
3 Schedule
Provide a chronological work plan with realistic milestones. Tips: Break tasks into concrete steps. Include checkpoints and buffer time. Use weekly or monthly units depending on project size. Seek advice from experienced researchers. Example Schedule Month Week Task Description 1 1 Planning Work plan, coordination 1 2–3 Literature Review Search and initial analysis 1 4 Literature Evaluation Critical review and selection 2 5–6 Theoretical Framework Develop theory and methodology 2 7 Concept Development Draft interview/ analysis guide 2 8 Scenario Planning Design simulations or experiments 3 9–10 Recruitment Acquire participants 3 11 Data Collection Conduct interviews/ experiments 3 12 Data Preparation Clean and pre-analyze data 4 13–14 Qualitative Analysis In-depth qualitative analysis 4 15–16 Quantitative Analysis Statistical analysis 5 17 Writing Results Compose results section 5 18 Discussion Interpret findings 5 19 Methodology Review Finalize methodology 5 20 Full Revision Edit and review all chapters 6 21 Final Editing Final proofreading, formatting 6 22–24 Finalization Buffer, printing, submission Reference [1] Helmut Balzert, Christian Schäfer, Marion Schröder, Uwe Kern, Roman Bendisch, and Klaus Zeppenfeld. Wissenschaftliches Arbeiten: Ethik, Inhalt und Form wissenschaftlicher Arbeiten, Handwerkszeug, Quellen, Projektmanagement, Präsentation. W3L-Verlag, Herdecke/Witten, 2011. ISBN: 978-3-937137-59-9.
Appendix C Additional Details on the Exposía Dataset
Exposés
Each exposé is available in two versions: a draft and a revised final version produced after receiving feedback.
Each version is provided as (i) a compiled PDF file and (ii) the original source, including the separate bibliography file (.bib) and all textual content.
Reviews
Every exposé draft is accompanied by two student reviews written by fellow students taking the course and at least one instructor review. Instructional staff included nine instructors and consisted of PhD students, postdoctoral researchers, and teaching assistants holding a Master’s degree (Appendix G.1 (Group 1)). The feedback was double-blind among students, but not with respect to the instructional staff. During the revision process, students could identify whether a specific feedback came from an expert or a student.
Each review is released as (i) an HTML document that preserves the original formatting used on the review platform and (ii) a plain-text rendering. For all reviews, we provide basic metadata including a synthetic reviewer name, reviewer role, document size, the associated exposé (if available), and an anonymized identifier.
For all reviews for which the author provided additional consent for behavioral logging, we release the full edit history of the review text as stored by the platform, capturing the successive revisions made by the reviewer while composing their feedback. This results in sequences of review drafts that can be aligned chronologically via timestamps.
Inline comments and voting
Beyond document-level feedback, the platform supports fine-grained, span-based comments on the exposé draft. For each exposé, we collect a set of inline comments created by students and instructors while reading the draft. Each comment is attached to a character-level span in the PDF-rendered text of the exposé and is associated with (i) optional free-text content, (ii) a categorical tag from a fixed taxonomy $\mathcal{Y}=\{\textit{Strength},\textit{Weakness},\textit{Highlight},\textit{Other}\}$ and (iii) metadata about the author role (student or instructor) and the underlying review document it belongs to.
The platform further allows other participants, but not the comment author, to cast binary votes (thumbs up / thumbs down) on each comment. These votes were intended as judgments of comment helpfulness. However, interviews with participants indicate that student authors sometimes used the voting mechanism as a personal checklist during revision (marking comments as done rather than as helpful), despite explicit instructions from the instructors to use it purely as a helpfulness signal. We therefore provide the raw voting data, together with information about voter role, but caution that votes do not form a clean ground truth for perceived usefulness.
Behavioral interaction logs
For users who provided explicit additional consent, we release interaction logs from the review platform, covering 45 students and 9 instructors. Each record contains an action type, a JSON-encoded data payload, a timestamp, and an anonymized user identifier. Logged events are timestamped UI interactions (Table 5); an overview of activity over the semester is shown in Figure 6 and discussed in Appendix L . Actions including navigation and visibility changes (e.g., Navigation step, PDF page show/hide, Browser Tab show/hide), scrolling and resizing (e.g., Annotation Page scroll, PDF scroll, Sidebar scroll, PDF (Browser) resize), interface controls and modal events (e.g., Sidebar click, Topbar click, Modal show/Modal hide, Modal Button click), and lightweight editing interactions (e.g., Text select, Copy text, Text paste).
Figure 6: Daily number of actions across the semester for instructors (n=9) and students (n=45). Bars show per-day activity (instructors in gray, students in blue). Alternating vertical shading indicates calendar weeks (Mon–Sun), and the hatched region marks the winter break. Dashed lines mark the start of reviews (green), the review deadline (red), and the revision deadline (orange). The y-axis is broken to accommodate the peak activity before the review deadline.
Action Students (n=45) Instructors (n=9) Total
Mouse move 73,811 131,272 205,083
PDF page show/hide 53,378 102,120 155,498
Annotation Page scroll 18,944 31,080 50,024
Browser Tab show/hide 16,810 31,657 48,467
PDF (Browser) resize 13,377 25,475 38,852
Text select 7,403 8,043 15,446
Sidebar scroll 2,445 7,248 9,693
Navigation step 2,705 3,742 6,447
Sidebar click 2,882 3,382 6,264
Topbar click 1,826 3,512 5,338
PDF scroll 2,206 2,179 4,385
Copy text 1,403 2,371 3,774
Table click 1,811 1,164 2,975
Text paste 1,208 1,435 2,643
Browser load 715 1,364 2,079
Modal show 889 324 1,213
Modal hide 396 304 700
Search used 102 369 471
Modal Button click 259 176 435
Total 202,570 357,217 559,787
Table 5: Number of behaviour data rows per action, split by user group.
Authorship metadata and anonymization
Each exposé, review, inline comment, vote, and interaction event is associated with an anonymized author identifier and a role label (student, junior, senior). To support more interpretable examples and qualitative analysis, we replace real names with consistently applied synthetic names generated with Faker. [^10] https://faker.readthedocs.io/en/master No raw platform identifiers, email addresses, or other personally identifying information are released. This setup enables the analysis of role-specific behavior and bias while preserving participant privacy.
Appendix D Criteria and Rubrics for Exposé and Review Assessment
This appendix provides an overview of exposé and review criteria used to grade the exposé and student reviews, respectively.
Expertise levels
Each criterion is assigned an expertise level [^11] Two co-authors inspected all criteria and labeled each for the expertise level required for scoring an exposé for a given criterion., indicating the type of judgment needed to assess the quality of the exposé or review with respect to this criterion. We introduce a three-level expertise scheme (expertise levels by criteria are shown in Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric for exposé and review criteria, respectively):
  • Level 0 (formal). No academic or expert knowledge required. Largely objective. Sample exposé criteria: Metadata available; Template used; Number of pages. Review criteria: None.
  • Level 1 (academic). Comprehension-based judgments of clarity, coherence, and basic research sense-making, but no specific in-domain expertise is required. Sample exposé criteria: Language quality; Hook existing. Sample review criteria: Summary available; Language.
  • Level 2 (expert). Expert judgment of the soundness and adequacy of the reasoning and its alignment with disciplinary standards. Sample exposé criteria: Teaser (RQ); Relevance of theoretical framework. Sample review criteria: Knowledge of correct response; Learning goals.
D.1 Exposé Criteria
Table LABEL:tab:expose_rubric lists all the 36 exposé assessment criteria, grouped into nine high-level rubrics. These criteria define the scoring rules applied to each submitted exposé.
Refinement of criteria and rubrics for exposé assessment
Before the course was run, the designed criteria and rubrics were piloted on exposés previously submitted within the department: two members of the instructional staff assessed together a small sample of old student exposés, and other instructional staff members went over these to give feedback. This feedback was used to verify that every criterion is assessable and the criteria definitions are clear. Criteria definitions were adjusted accordingly for clarity and consistency.
D.2 Review Criteria
Table LABEL:tab:review_rubric lists all the 24 review assessment criteria grouped into eight high-level rubrics. These criteria define the scoring rules applied to each submitted student review.
Refinement of criteria and rubric for review assessment
As with the exposé rubrics, all criteria and rubrics were reviewed by instructors and piloted on sample reviews before course deployment.
Table 6: 36 criteria used to grade student exposés. We organize the criteria into nine rubrics, by grouping together criteria that evaluate related aspects of exposé writing.
Code Criterion Expert Point descriptors
level
Rubric: Language/Quality (LA)
Assessment of language quality and structural coherence of the text.
LA1 Language quality 1 0: many linguistic errors (e.g. wrong tense, wrong personal pronouns)
1: few linguistic errors, clear and comprehensible sentences
2: no linguistic errors, clear and comprehensible sentences with consistent terminology
LA2 Common Thread 1 0: No recognizable common thread, the structure is unclear and jumpy
1: Partly recognizable, but the structure is not continuous
2: Text has a continuous structure that is easy to follow, common thread is clear and recognizable from beginning to end
Rubric: Metadata (ME)
Evaluation of title appropriateness and creativity.
ME1 Preliminary title 1 0: Title does not match the topic
1: Title fits the topic; similar title to the one in the topic suggestions
2: Title fits the topic; creative title changed in a novel way and/or fits into the story of the exposé
ME2 Metadata available 0 0: Name, date and matriculation number are not entered correctly/changed
1: Name, date, matriculation number are correct and updated
Rubric: Form/Structure (ST)
Evaluation of document formatting and structural requirements.
ST1 Template used 0 0: Latex template is not used
1: Latex template is used
ST2 Number of pages 0 0: > three pages (motivation + approach only)
1: <= three sides (motivation + approach only)
Rubric: Motivation (MO)
Assessment of problem identification, context establishment, and research question formulation.
MO1 Hook existing 1 0: Hook not available
1: Hook present and suitable for the topic
MO2 Anker 1 0: The problem is not mentioned at all.
1: The problem is stated in general terms without specific details. It is recognizable which problem is being addressed, but important information is omitted.
2: The problem is described in detail. All relevant aspects of the problem are mentioned and clarify its scope and possible effects.
MO3 Domain 1 0: The domain or context in which the problem occurs is not mentioned.
1: The domain or the relevant expert area of the problem is mentioned. It is clear in which environment the problem exists.
MO4 Problem relevance 1 0: The relevance or importance of the problem is not addressed. It remains unclear why this problem is relevant or worth discussing.
1: The importance and relevance of the problem is described. It is made clear why the problem is important and what impact or consequences it could have (e.g. for specific individuals, organisations or society as a whole).
MO5 Problem handling 1 0: current handling of the problem or current criticism is not given
1: current handling of the problem or current criticism is given
MO6 Teaser (RQ) 2 0: There is no research question available
1: A research question exists, but it has weaknesses, e.g. due to unclear terminology or vague formulation. It is difficult to recognise what is to be investigated and how the question fits the problem description.
2: The research question is formulated clearly and precisely. It clearly conveys what is to be investigated, and fits thematically with the problem description.
MO7 RQ Limitation 2 0: The research question is either too general or too broad, so that it does not seem realistic to answer it within the scope of a Bachelor’s thesis, to answer it in the context of a Bachelor’s thesis.
1: The research question is precise and clearly defined so that it can realistically be addressed in the context of a Bachelor’s thesis. The question is formulated in concrete terms and describes specifically which aspects are to be investigated without deviating from the topic.
Rubric: Approach (AP)
Assessment of state of the art analysis, theoretical framework.
AP1 SOTA 1 0: SOTA not available
1: SOTA present and correctly cited
AP2 SOTA Relevance 1 0: The relevance of SOTA is not present or unclear.
1: The relevance of SOTA is mentioned, but is only partially comprehensible or weakly presented.
2: It is clearly shown how SOTA is relevant to the topic of the exposé and why it is relevant to the research.
AP3 SOTA Weaknesses 1 0: SOTA’s weak points are not mentioned
1: SOTA’s weaknesses are mentioned, but only superficially or unclearly described
2: SOTA’s weaknesses are clearly identified and precisely explained; it is clear which weaknesses are relevant for own research
AP4 SOTA Delimitation 1 0: No differentiation of own performance from the current state of research
1: The own achievement is differentiated from SOTA; it is explicitly emphasized, what is new or unique about your own research
AP5 SOTA Combination 2 0: No meaningful combination of existing material (even if reference is made to existing material, but its combination or consolidation is unconvincing or unclear)
1: Existing material is clearly and meaningfully combined; the added value of the combination is clear.
AP6 Theoretical Framework 1 0: The theoretical framework does not exist or is completely misleading
1: The theoretical framework is present, but unclear, poorly structured or difficult to understand
2: The theoretical framework is clearly and logically structured. The theoretical approaches are presented in a understandable and comprehensible
AP7 Relevance of theoretical framework 2 0: The theoretical framework shows no connection to the research topic, or the theory is irrelevant
1: The theoretical framework clearly demonstrates how the selected theories directly relate to the research topic and support the research question
AP8 Methodology Availability 1 0: Methodology not available
1: Methodology available and suitable for answering the RQ(s)
Rubric: Methodology (MD)
Assessment of the methodology.
MD1 Methodology Completeness 1 0: Methodology is available, but not all RQ(s) are addressed
1: Methodology is available and addresses all points of the research questions
MD2 Methodology Relevance 1 0: The relevance of the methodology for the research project is not clear or is unclear
1: It is clearly and precisely demonstrated why the chosen methodology is relevant to the research project and how it contributes to the research questions
MD3 Methodology Target group 1 0: The target group is not specified or there is no precise differentiation
1: The target group is clearly and precisely defined and it is clear why it is relevant to the methodology
MD4 Methodology Existing Material 1 0: No reference is made to the existing material or the material is inappropriate
1: The existing material is appropriately and accurately related to the topic; it is clear that this material is relevant to the methodology
MD5 Methodology Difficulties 2 0: Potential difficulties in the practical implementation of the methods are not addressed. Potential challenges that could hinder the research process are overlooked.
1: Potential difficulties in the practical implementation of the methods are described clearly and precisely. Concrete solutions or strategies to overcome these difficulties are also presented, in order to organize the research process as efficiently and target-oriented as possible.
MD6 Methodology Possibilities/restrictions 2 0: The possibilities or limitations of the methods used are not addressed, or they are only mentioned superficially, without concrete details. It remains unclear how these methods contribute to answering the research question and which limitations could possibly influence the validity of the results.
1: The possibilities and limitations of the methods are described clearly and precisely. It is clear what strengths the methods bring to the study and what weaknesses or limitations could possibly influence the results. limitations could possibly influence the results. The choice of methods is critically reflected upon and related to the research project.
MD7 Methodology Details 2 0: No additional explanations are given on the methodology or implementation
1: The methodology is explained comprehensively and precisely with additional explanations (e.g. details on implementation), so that the procedure is clearly comprehensible
Rubric: Schedule (SC)
Assessment of project timeline structure, completeness, and realism.
SC1 Schedule Availability 0 0: Tabular schedule not available
1: Tabular schedule in chronological order available
SC2 Schedule Completeness 1 0: Tabular schedule contains gaps
1: Tabular schedule available from start to finish and reasonably divided into blocks
SC3 Schedule Block Description 0 0: not available or only headlines available
1: individual blocks described (not only headlines)
SC4 Schedule Realistic Relevance 1 0: The time distribution is obviously unrealistic and important work steps are missing. The research questions are not or only insufficiently addressed
1: The timeline contains the most important steps, but some of the timelines are unrealistic or unclear. It is partially unclear how the steps contribute to answering the research questions. The timeline does not appear to be fully aligned with the Bachelor’s thesis context.
2: The timeline is clearly structured, realistic and aligned with the scope of the Bachelor’s thesis. All important work steps are included and organised in a sensible time frame. Each step contributes to answering the research questions, and the entire plan fits coherently into the exposé and the context of the thesis.
Rubric: Bibliography (BI)
Assessment of bibliography consistency and literature relevance.
BI1 Bibliography Consistency 0 0: Bibliography with different citation styles
1: Bibliography is standardized ( consistent citation style) and essentially all information on the literature is available
BI2 Key literature 2 0: Literature does not fit the topic
1: Literature fits the topic
2: Literature fits the topic and most of it is highly relevant
Rubric: Additional points (AD)
Allow additional points for very good submissions and major mistakes in the submission.
AD1 Additional points 1 0: No additional strengths beyond existing criteria.
1: Some additional strengths beyond existing criteria.
2: Clear exceptional strengths beyond existing criteria.
AD2 Negative points 1 0: No major issues beyond existing criteria.
-1: Notable issues beyond existing criteria.
-2: Serious issues that strongly reduce overall quality.
Table 7: 24 criteria used to grade student reviews. We organize the criteria into eight rubrics, by grouping together criteria that evaluate related aspects of review.
Code Criterion Expert Point descriptors
Rubric: Structure and Clarity (SC)
Assessment of overall structure and clarity.
SC1 Summary available 1 0: No summary available or, the summary does not reflect understanding of the performance
1: Simple summary that demonstrates basic understanding of the performance
2: Concise and accurate summary that clearly reflects understanding of performance and builds confidence in feedback
SC2 Clarity 2 0: Feedback needs considerable improvement in clarity and comprehensibility, comments are disjointed or difficult to follow
1: Feedback is largely clear, but sometimes lacks specific references to particular lines, pages, sections or figures/tables
2: Feedback is clear, well-structured and easy to understand, with coherent comments and precise references to specific points in the text
Rubric: Language (LG)
Assessment of language quality and tone.
LG1 Language 1 0: Many linguistic errors (e.g. wrong tense, wrong personal pronouns)
1: Almost no linguistic errors, clear and comprehensible sentences with standardised terminology
LG2 Tone 1 0: The tone is inappropriate, sounds judgemental or reproachful, criticism comes across as destructive or personal
1: The tone is generally respectful, avoids personal attacks, but remains unclear or less friendly in places
2: The tone is consistently calm, respectful and polite; criticism is constructive, factual and only directed at the text, not the person
Rubric: Feed Up (FU)
Assessment of learning goals and success criteria.
FU1 Learning Goals 2 0: No learning objectives, unclear what is to be achieved, no orientation
1: Partly unclear learning objectives, generally formulated, offer only limited orientation
2: Clear learning objectives, specific, challenging, comparable and provide clear direction
FU2 Success Criteria 1 0: No success criteria, success unclear or vaguely formulated, objectives not defined
1: Partly unclear success criteria, available but imprecise or general, orientation partly missing
2: Clear success criteria, concrete, measurable goals and clear definition of success
Rubric: Feed Back (FB)
Assessment of feedback quality and knowledge transfer.
FB1 Knowledge of result 1 0: No information as to whether the task has been solved or not
1: Information given as to whether the task has been solved or not
FB2 Knowledge of correct response 2 0: No information on the correct answer
1: Correct answer is clearly communicated
FB3 Knowledge about task constraints 1 0: No information on rules, requirements or restrictions of the task
1: Clear information on rules, requirements or restrictions of the task (e.g. specifications)
FB4 Knowledge about concepts 1 0: No information on conceptual knowledge
1: Basic information on relevant concepts is provided
FB5 Knowledge about mistakes 1 0: No information about errors
1: Information on the number, location, source or type of errors
FB6 Self-feedback 1 0: Self-feedback is available that has not been combined with task, process or self-regulation (including feedback on the learner’s abilities (i.e. intelligence or talent))
1: No self-feedback is available or, if available, in combination with task, process or self-regulation
Rubric: Feed Forward (FF)
Assessment of future learning guidance.
FF1 Self-regulation 1 0: No indication of how to approach the task, no recognisable approaches to self-monitoring or self-assessment
1: Either indications of strategies for working on the task and recognising errors (e.g. suggestions for improvement, targeted search for information) or approaches to self-monitoring and self-assessment (e.g. reflection on own understanding, strategies or willingness to seek help)
FF2 Learning Skills 1 0: No information on how to develop or improve learning skills
1: Information on how to improve the learning objective and how to deal with challenges
Rubric: Content Quality (CQ)
Assessment of content accuracy and usefulness.
CQ1 Correctness 2 0: The content has significant deficiencies in terms of correctness
1: The content is largely correct
2: The content is completely correct
CQ2 Actionability 1 0: No specific instructions available, feedback is difficult to implement
1: Some instructions available, but only partially clear or specific
2: Clear, specific feedback with actionable instructions, e.g. comparisons with existing methods, application to other tasks, definitions, explanations, discussions, additional analyses or specific suggestions to remove confusion
CQ3 Argumentation 1 0: Statements or conclusions are not supported by evidence or comprehensible explanations. The argumentation remains superficial or unsystematic, making the content unconvincing.
1: Statements are supported by comprehensible evidence, examples or explanations. The argumentation shows a clear link between evidence and conclusions, which strengthens the quality of the content.
Rubric: Errors (ER)
Assessment of potential errors and issues in feedback.
ER1 Neglect 2 0: All important details have been considered and no unnecessary questions are asked
-1: Important details have been overlooked, leading to unnecessary questions or criticism
ER2 Vague Critic 1 0: Criticism is specific and clearly states what is missing or can be improved
-1: Unspecific criticism, mentions missing components without clearly stating what is missing
ER3 Out-of-Scope 1 0: Proposals remain within the scope of the task and are relevant
-1: Suggestions refer to methods or analyses that go beyond the intended scope
ER4 Missing Reference 1 0: Alternative methods, etc. (if available) are supported with justification or references
-1: Alternative methods, etc. (if available) are suggested without justification or references
ER5 Contradiction 1 0: Feedback is consistent in itself, no contradictory statements
-1: Contradictory feedback, e.g. criticism and praise for the same methods
Rubric: Additional points (AD)
Allow additional points for very good reviews or deduct points for major mistakes.
AD1 Additional points 1 0: No additional strengths identified.
1: Notable additional strengths beyond other criteria.
AD2 Negative points 1 0: No additional major issues identified.
-1: Significant issues beyond other criteria.
Appendix E Summary of Collected Sources for Exposé Criteria
Source Lang BA MA Diss. Grant/Proj. Link
TU Berlin EN [1]
FU Berlin EN [2]
Uni Leipzig EN [3]
Uni Magdeburg EN [4]
Uni Freiburg DE/EN [5]
Hochschule Magdeburg EN [6]
Uni Hannover EN [7]
TU Dresden EN [8]
Hochschule für Politik München EN [9]
LMU München EN [10]
Uni Wien EN [11]
Uni Kassel EN [12]
Uni Mannheim EN [13]
FH Erfurt EN [14]
Scribbr (DE) DE [15]
WU Wien DE [16]
Finito24 DE [17]
Uni Bremen EN [18]
Table 8: University and institutional exposé guidelines with explicit scope for Bachelor, Master, Dissertation, or Grant/Project proposals.
Across the full corpus of 110 exposé guidelines we reviewed, we identified materials from universities and higher-education institutions in Germany, Austria, and Switzerland, complemented by student-facing writing portals (e.g., Scribbr, Finito24, Studieren.de), which we further investigated. In terms of disciplinary scope, our primary sources are published by faculties and units in the humanities and education, the social sciences (including sociology and political science), medicine, economics and business, as well as applied/STEM fields (i.e., architecture, computer science, and civil engineering). Roughly three-quarters of the sources are published in English, primarily from internationalized programs (e.g., TU Berlin, Universität Wien, TU Dresden, Uni Magdeburg), while the remainder appear in German and serve national undergraduate or graduate audiences. About one-fifth of the documents explicitly target Bachelor- or Master-level theses, and a smaller subset focuses on doctoral or dissertation exposés, such as those from the University of Leipzig Graduate Centre and the University of Magdeburg Medical Faculty. Only a few pages, most notably the TU Berlin departmental guide, explicitly include project or grant proposal contexts.
Thematically, nearly all guidelines converge on the same structural expectations: a clear Motivation or problem statement, a theoretically informed Approach or methodology, a feasible Schedule or timeline, and a complete Bibliography. Additional sections, such as Research Questions, Related Work, and Formal Aspects (including format, citation, or layout), occur frequently, especially in English-language handbooks and institutional templates.
Across this heterogeneity, the documents show remarkable consistency in pedagogical intent: to ensure students articulate a focused question, justify methodological choices, and plan realistic execution. These recurrent patterns directly informed the rubric domains of Motivation, Approach, Schedule, and Bibliography used for annotation and model benchmarking in our dataset.
Links for Table 8:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
Appendix F Exposía Dataset Statistics
This section provides additional descriptive statistics for Exposía that complement the summary in Section 5 of the main paper. To give a more detailed view of the dataset’s composition, we report basic statistics on the exposés, reviews, and assessment scores.
F.1 Exposé Revision Statistics
Table 9 reports section-wise text statistics for Motivation, Approach, and Schedule in draft vs. final versions (mean $\pm$ SD). Raw is converted to plain text with pylatexenc [^12] https://github.com/phfaist/pylatexenc; sentence/word segmentation uses spaCy. We report length/structure (#Words, #Sentences and average word length), lexical/readability (type–token ratio (TTR) and Flesch Reading Ease (FRE); Flesch, 1948 via textstat [^13] https://github.com/textstat/textstat), and draft–final similarity (cosine similarity of SentenceTransformer embeddings141414sentence-transformers/all-mpnet-base-v2 and normalized token-level Levenshtein similarity; Levenshtein, 1966).
Across sections, final exposés are longer in words and sentences, with the largest expansion in Approach; average word length changes only marginally. Lexical diversity (TTR) decreases slightly from draft to final in all three sections. The negative FRE values for Schedule likely reflect formatting artifacts (fragmentary, table-like content) interacting with plain-text conversion and segmentation.
Similarity and topic structure
High embedding cosine similarities in Table 9 indicate strong semantic continuity between draft and final, even where token-level edit similarity is lower (notably for Motivation and Approach). Table 10 reports pairwise embedding cosine similarities within and across topics: within-topic pairs are generally more similar than cross-topic pairs, with the largest within-topic variance for Generative AI in Healthcare.
Motivation Approach Schedule
Metric Draft Final Draft Final Draft Final
Length and structure
#Words 402 ($\pm$133) 428 ($\pm$125) 748 ($\pm$258) 851 ($\pm$221) 248 ($\pm$83) 270 ($\pm$131)
#Sentences 21.8 ($\pm$8.5) 24.0 ($\pm$7.8) 40.9 ($\pm$18.4) 47.7 ($\pm$15.3) 7.3 ($\pm$8.3) 8.4 ($\pm$8.7)
Avg. word length 5.38 ($\pm$0.45) 5.43 ($\pm$0.44) 5.41 ($\pm$0.49) 5.46 ($\pm$0.50) 5.55 ($\pm$0.37) 5.50 ($\pm$0.36)
Lexical and readability
Lexical diversity 0.54 ($\pm$0.06) 0.53 ($\pm$0.06) 0.45 ($\pm$0.07) 0.43 ($\pm$0.06) 0.60 ($\pm$0.06) 0.59 ($\pm$0.06)
Flesch reading ease 43.0 ($\pm$14.5) 41.6 ($\pm$13.3) 42.1 ($\pm$15.3) 41.0 ($\pm$15.9) -50.4 ($\pm$90.4) -34.2 ($\pm$90.1)
Draft–Final similarity
Cosine similarity 0.96 ($\pm$0.04) 0.95 ($\pm$0.06) 0.93 ($\pm$0.14)
Levenshtein similarity 0.64 ($\pm$0.24) 0.62 ($\pm$0.26) 0.81 ($\pm$0.24)
Table 9: Chapter-level textual statistics for exposés by version (draft vs. final). Entries show mean (±SD) for length, lexical diversity, readability, and draft–final similarity metrics.
Motivation Approach Schedule
Topic Draft Final Draft Final Draft Final
T1 0.59 ($\pm$0.13) 0.58 ($\pm$0.13) 0.55 ($\pm$0.14) 0.56 ($\pm$0.14) 0.52 ($\pm$0.16) 0.52 ($\pm$0.16)
T2 0.60 ($\pm$0.12) 0.66 ($\pm$0.09) 0.60 ($\pm$0.11) 0.62 ($\pm$0.12) 0.59 ($\pm$0.13) 0.61 ($\pm$0.12)
T3 0.48 ($\pm$0.14) 0.47 ($\pm$0.14) 0.40 ($\pm$0.17) 0.41 ($\pm$0.16) 0.54 ($\pm$0.13) 0.50 ($\pm$0.11)
Overall 0.37 ($\pm$0.18) 0.38 ($\pm$0.18) 0.35 ($\pm$0.18) 0.35 ($\pm$0.19) 0.48 ($\pm$0.16) 0.46 ($\pm$0.15)
Table 10: Pairwise cosine similarity between exposés (each exposé is compared with all others within the same topic and version). Entries show mean (±SD) over all unordered document pairs, computed separately per chapter. Topic mapping: T1=The Holographic AI Assistant, T2=Interactive LLM-based teaching agents, T3=Generative AI in Healthcare.
Bibliography size and stability
Table 11 reports topic-wise bibliography statistics for draft vs. final exposés (mean $\pm$ SD), measured as the number of distinct Bib entry keys and their draft–final overlap (Jaccard similarity over key sets). Draft bibliography size differs substantially by topic (T1 << T2 << T3). Despite these initial gaps, revisions show a consistent pattern across topics: authors add about two references on average while removing fewer than one, suggesting modest expansion rather than wholesale replacement. Correspondingly, mean draft–final Jaccard similarity remains relatively high (overall $0.78$), indicating that bibliographies are generally stable across revisions with targeted additions.
Draft Final Similarity
Topic $\bar{n}_{\text{draft}}$ $\bar{n}_{\text{final}}$ $\bar{n}_{\text{added}}$ $\bar{n}_{\text{removed}}$ $\overline{\text{Jaccard}}$
T1 $6.74\,\pm\,6.02$ $8.63\,\pm\,5.63$ $2.16\,\pm\,3.86$ $0.26\,\pm\,0.56$ $0.73\,\pm\,0.40$
T2 $9.22\,\pm\,6.39$ $10.78\,\pm\,6.31$ $2.06\,\pm\,2.62$ 0.50± $50\,\pm\,1.2$5 $0.76\,\pm\,0.27$
T3 $11.83\,\pm\,8.01$ $13.06\,\pm\,8.74$ $1.89\,\pm\,2.97$ $0.67\,\pm\,2.83$ $0.84\,\pm\,0.26$
Overall $9.22\,\pm\,7.04$ $10.78\,\pm\,7.11$ $2.04\,\pm\,3.15$ $0.47\,\pm\,1.77$ $0.78\,\pm\,0.32$
Table 11: Bibliography size and draft–final changes by topic. We report mean (±SD) number of entries in draft and final versions, average additions/removals, and mean draft–final entry-set Jaccard similarity computed over Bib keys. Topic mapping: T1=The Holographic AI Assistant, T2=Interactive LLM-based Teaching Agents, T3=Generative AI in Healthcare.
Draft–final improvements on rubric scores
Table 12 reports per-rubric statistics for draft and final exposés. It complements Section 5.1 by additionally showing (i) the remaining gap between the mean final score and the rubric maximum and (ii) the observed score ranges (min–max) for each version. As in the main text (Section 5.1), the largest mean gains occur in conceptually central rubrics such as Approach and Motivation. In contrast, smaller changes in Metadata, Bibliography, and Form/Structure largely reflect limited headroom: these rubrics have low maxima, and draft scores are already close to the ceiling. Finally, Additional points changes little because it represents bounded instructor adjustments rather than points that students can systematically earn through revision (see Appendix F.3 for the corresponding scoring rules). Overall, the revision signal is therefore most visible in rubrics with substantial remaining headroom, and is concentrated in substantive content and argumentation.
Mean (±SD) Gaps Observed bounds
Rubric Draft Final $\Delta$ Gap Max Draft Final
Approach 7.91 ($\pm$2.25) 9.40 ($\pm$2.06) 1.49 ($\pm$1.74) 1.60 11.00 1.00–11.00 1.00–11.00
Motivation 6.93 ($\pm$1.46) 8.09 ($\pm$1.01) 1.16 ($\pm$1.17) 0.91 9.00 4.00–9.00 6.00–9.00
Language/Quality 2.69 ($\pm$1.02) 3.45 ($\pm$0.78) 0.76 ($\pm$0.81) 0.55 4.00 0.00–4.00 1.00–4.00
Methodology 4.07 ($\pm$1.32) 4.76 ($\pm$0.57) 0.69 ($\pm$1.08) 0.24 5.00 0.00–5.00 3.00–5.00
Schedule 4.15 ($\pm$0.84) 4.65 ($\pm$0.55) 0.51 ($\pm$0.83) 0.35 5.00 0.00–5.00 3.00–5.00
Bibliography 2.56 ($\pm$0.85) 2.91 ($\pm$0.29) 0.35 ($\pm$0.77) 0.09 3.00 0.00–3.00 2.00–3.00
Metadata 2.75 ($\pm$0.51) 2.95 ($\pm$0.23) 0.20 ($\pm$0.48) 0.05 3.00 1.00–3.00 2.00–3.00
Additional points -0.02 ($\pm$0.36) 0.00 ($\pm$0.27) 0.02 ($\pm$0.30) 2.00 2.00 -2.00–1.00 -1.00–1.00
Form/Structure 1.96 ($\pm$0.19) 1.96 ($\pm$0.19) 0.00 ($\pm$0.19) 0.04 2.00 1.00–2.00 1.00–2.00
Table 12: Per-rubric statistics for draft and final exposé scores. We report mean (±SD), average remaining gap to the rubric maximum, and observed score bounds (min–max) for each version.
F.2 Review Statistics
Table 13 reports text statistics for review documents by author role (mean $\pm$ SD), computed analogously to the exposé statistics. Student reviews are substantially longer than instructor reviews, while instructors show higher lexical diversity.
Review author role
Metric Student Junior Instructor Senior Instructor
Length and structure
#Words 680 ($\pm$455) 285 ($\pm$134) 347 ($\pm$116)
#Sentences 33.7 ($\pm$21.3) 16.7 ($\pm$8.4) 18.7 ($\pm$6.3)
Avg. word length 5.31 ($\pm$0.44) 5.16 ($\pm$0.28) 5.18 ($\pm$0.49)
Lexical and readability
Lexical diversity 0.46 ($\pm$0.08) 0.56 ($\pm$0.09) 0.55 ($\pm$0.08)
Flesch reading ease 50.2 ($\pm$13.3) 58.9 ($\pm$9.9) 54.3 ($\pm$14.2)
Table 13: Textual statistics of review texts by author role. Entries show mean ($\pm$ SD).
Instructor roles
We distinguish between junior and senior instructors based on their academic stage and experience in reviewing and research. Junior instructors are recruited Master’s students who have completed a Bachelor’s thesis and thus have basic academic writing experience, whereas senior instructors are course staff (PhD researchers and a postdoc) with substantially more experience in academic peer review and supervision; see Appendix G.1 for details on rater composition and training.
Inline comment distributions
Table 14 summarizes comment volume and length by role: junior instructors contribute the most comments overall, while students and senior instructors write longer comments on average. Table 15 reports the distribution of span-level tags. Totals are similar but not identical across the two tables because span annotations are only recorded for comments that are attached to a specific text span; page-level comments have no associated span annotation. Across roles, Weakness is the most frequent tag, followed by Strength, Other, and Highlight.
Role Comments Avg chars Avg words
St. (n=31) 640 77.0 12.6
Jr. (n=7) 1121 53.8 9.1
Sr. (n=5) 492 68.8 12.0
$\Sigma$ (n=43) 2253 63.6 10.7
Table 14: Basic statistics of comments by role, including total counts and average length in characters and words. St.=Students, Jr.=Junior Instructors, Sr.=Senior Instructors.
Tag Metric St. (N=31) Jr. (N=7) Sr. (N=5) $\Sigma$
Highlight Raw count 90 28 58 176
Avg. length 112.68 ($\pm$188.62) 122.79 ($\pm$149.17) 168.34 ($\pm$227.29) 132.63 ($\pm$198.48)
Other Raw count 107 145 52 304
Avg. length 120.66 ($\pm$236.88) 41.77 ($\pm$61.91) 90.83 ($\pm$132.55) 77.93 ($\pm$160.87)
Strength Raw count 131 200 73 404
Avg. length 254.63 ($\pm$263.11) 178.40 ($\pm$239.87) 196.52 ($\pm$285.21) 206.39 ($\pm$258.44)
Weakness Raw count 308 691 290 1289
Avg. length 79.20 ($\pm$106.17) 59.17 ($\pm$127.02) 97.88 ($\pm$154.42) 72.66 ($\pm$130.22)
$\Sigma$ Raw count 636 1064 473 2173
Table 15: Distribution of span-level annotation tags by role. We report raw counts and average span length (±SD). St.=Students, Jr.=Junior Instructors, Sr.=Senior Instructors.
F.3 Assessment scores coverage
Table 16 summarizes the number of assessments that include a complete set of criterion-based scores for exposé drafts and final submissions, by rater group. Draft exposés were graded by instructors during the semester (Group 1) and re-scored after the course by Group 2 to compute inter-annotator agreement (Section 5.2). Final exposés, in contrast, were not re-scored post-semester: due to the substantial time required for criterion-based grading, Group 2 focused on a draft subset rather than duplicating the full set of final-submission grades. A similar pattern holds for student review grading (Table 17): reviews were graded during the semester and selectively re-scored for agreement analyses, but criterion-based scores are only available for student-written reviews (not for instructor-written reviews).
Group 1 Group 2
Version / Topic Jr. Sr. $\Sigma$ Jr. Sr. $\Sigma$
Draft (N = 55)
T1 (N=19) 14 5 19 2 17 19
T2 (N=18) 14 4 18 4 14 18
T3 (N=18) 15 3 18 2 16 18
$\Sigma$ Draft 43 12 55 8 47 55
Final (N = 55)
T1 (N=19) 15 4 19 0 0 0
T2 (N=18) 14 4 18 0 0 0
T3 (N=18) 15 3 18 0 0 0
$\Sigma$ Final 44 11 55 0 0 0
Table 16: Distribution of criterion-based exposé assessments across topics (T1–T3), rater roles (Jr. Junior Instructors, Sr. Senior Instructors), rater groups (Group 1 in-semester, Group 2 post-semester), and versions (draft vs. final). T1: “The Holographic AI Assistant”; T2: “Interactive LLM-based teaching agents”; T3: “Generative AI in Healthcare”.
Group 1 Group 2
Reviews Jr. Sr. $\Sigma$ Jr. Sr. $\Sigma$
Student reviews (N = 110)
$E$ (N = 34) 28 6 34 0 59 59
$\neg E$ (N = 76) 67 9 76 0 111 111
$\Sigma$ 95 15 110 0 170 170
Table 17: Distribution of criterion-based student review assessments by instructor role (Jr. Junior Instructors, Sr. Senior Instructors) and rater group (Group 1 vs. Group 2), stratified by whether the reviewed author has an associated exposé ($E$) or not ($\neg E$).
Annotation effort
We emphasize that constructing this dataset is unusually labor-intensive. In Group 1, producing a full review with inline comments and criterion-based scores typically required on the order of 2–4 hours early in the semester, decreasing to roughly 1–2 hours as raters gained experience. Criterion-based grading itself is also high effort: grading a single exposé typically takes about 1–2 hours, while grading a single review requires 20–40 minutes. Despite these restrictions, the dataset currently comprises 55 exposés and 306 reviews and includes unusually rich annotation artifacts: inline comments and comment tags, free-text reviews, draft and final exposé versions, and hundreds of instructor assessment scores across 36 exposé and 24 review criteria, which we use throughout our analyses.
Scoring and special rubric behavior
Within each submission (draft exposé, final exposé, peer feedback), criterion-level scores are summed to yield rubric points, and rubric points are summed to obtain total submission points. Most rubrics use simple addition; we introduce a small number of rubric-specific rules to maintain consistency across topics and feedback situations. The exposé rubric Methodology is capped at 5 points, and the review rubric Feedback is capped at 4 points. For example, in the exposé rubric, Existing Material may be inapplicable when the proposed methodology relies on newly collected primary data rather than reusing an existing dataset; the cap ensures that full credit remains attainable by satisfying an appropriate subset of criteria. The review rubric Errors employs a deduction-only scheme, starting at 4 points and decreasing by 1 point for each identified error, with a minimum score of 0. Finally, the instructor-only rubric Additional points permits bounded adjustments in exceptional cases (Exposé: [2,+2][-2,+2], Review: [1,+1][-1,+1]); these adjustments do not increase the maximum attainable points.
Group 1 Group 2
Reviews per exposé N Jr. Sr. St. $\Sigma$ N Jr. Sr. $\Sigma$
Exposés (total N = 55)
0 review(s) 0 0 0 0 0 30 0 0 0
1 review(s) 23 17 6 0 23 24 4 20 24
2 review(s) 23 21 7 18 46 1 0 2 2
3 review(s) 9 8 3 16 27 0 0 0 0
$\Sigma$ totals 55 46 16 34 96 55 4 22 26
Orphaned reviews (total N = 184)
$\Sigma$ totals 184 88 20 76 184 0 0 0 0
Table 18: Distribution of reviews per exposé by review group (Group 1 vs. Group 2). In the Exposés block, N is the number of exposés in the bucket; role columns (Jr./Sr./St.) are counts of reviews written by that role (columns that are all-zero for a group are omitted); and $\Sigma$ is the total number of reviews. Orphaned reviews are reported separately.
Appendix G Inter-annotator agreement
This section documents the computation of inter-annotator agreement (IAA), complementing the criterion-based scoring description in Section 5.2 for Exposía. It also provides the reference point for interpreting LLM–human agreement in the main experiments. Because our criterion scores are discrete and ordinal, we prioritize agreement measures that account for ordered categories and report distributional diagnostics for criteria affected by prevalence skew.
G.1 Raters, training, and scoring design
Raters and training
Twelve raters participated and were organized into two groups. In total, 7 were junior instructors (student helpers) and 5 were senior instructors (4 PhD researchers and 1 postdoc). One PhD researcher participated in both groups.
Group 1 (in-semester): 6 student helpers, 2 PhD researchers, 1 postdoc; weekly calibration (jointly graded initial exposés, then addressed edge cases). Because grades were at stake, Group 1 was instructed to resolve uncertainty in favor of the student.
Group 2 (post-semester): 1 student helper, 3 PhD researchers; scored subsets of draft exposés and reviews using a recorded calibration video prepared by a Group-1 senior instructor.
Rater groups and coverage
Table 16 indicates that Group 1 grades are predominantly assigned by junior instructors, whereas Group 2 grades are predominantly assigned by senior instructors. This shift mainly reflects post-semester availability: once the course ended, much of the teaching staff was no longer available, so re-grading was carried out primarily by senior instructors; final submissions were graded only during the semester.
Scope of IAA
Our agreement analyses are necessarily restricted to items (exposés and reviews) that have assessments from two instructors. For exposés, we report agreement only for draft submissions, as final submissions were not scored post-semester due to the substantial time required for criterion-based assessment (Appendix F.3). For reviews, we focus on the subset that has criterion-based grades from Group 1 (student reviews). Finally, for a subset of reviews ($n=60$), we collected duplicate retrospective ratings from Group 2, enabling a within-group agreement comparison in addition to cross-group agreement (Table 21). Beyond criterion-level results, we also report aggregate agreement (i) by rubric domain and (ii) by criterion expertise level (see Appendix D).
Code Criterion Distribution (%) QWA $\alpha$ $r$ $\Delta\mu$
LA1 Language quality 0 (15), 1 (55), 2 (30) 0.82 0.15 0.17 0.13
LA2 Common Thread 0 (4), 1 (65), 2 (32) 0.87 0.05 0.20 0.38
Average Language 0.85 0.10 0.19 0.25
ME1 Preliminary title 0 (3), 1 (42), 2 (55) 0.86 0.10 0.31 0.44
ME2 Metadata available 0 (10), 1 (90) 0.80 -0.10 0.20
Average Metadata 0.83 0.00 0.31 0.32
ST1 Template used 0 (2), 1 (98) 0.96 -0.01 -0.02 0.00
ST2 Number of pages 0 (3), 1 (97) 0.98 0.66 0.70 0.02
Average Form/Structure 0.97 0.33 0.34 0.01
MO1 Hook existing 0 (15), 1 (85) 0.76 0.10 0.10 0.05
MO2 Anker 0 (5), 1 (63), 2 (33) 0.85 -0.00 0.06 0.31
MO3 Domain 0 (5), 1 (95) 0.95 0.38 0.49 0.05
MO4 Problem relevance 0 (17), 1 (83) 0.65 -0.20 -0.21 -0.02
MO5 Problem handling 0 (35), 1 (65) 0.64 0.20 0.24 0.18
MO6 Teaser (RQ) 0 (7), 1 (56), 2 (36) 0.85 0.15 0.21 0.22
MO7 RQ Limitation 0 (46), 1 (54) 0.58 0.17 0.25 0.27
Average Motivation 0.76 0.11 0.16 0.15
AP1 SOTA 0 (13), 1 (87) 0.85 0.35 0.37 0.07
AP2 SOTA Relevance 0 (9), 1 (56), 2 (35) 0.86 0.28 0.39 0.36
AP3 SOTA Weaknesses 0 (23), 1 (55), 2 (22) 0.83 0.25 0.30 0.27
AP4 SOTA Delimitation 0 (30), 1 (70) 0.73 0.36 0.40 0.16
AP5 SOTA Combination 0 (34), 1 (66) 0.65 0.23 0.29 0.20
AP6 Theoretical Framework 0 (20), 1 (47), 2 (33) 0.78 0.15 0.20 0.29
AP7 Relevance of theoretical framework 0 (35), 1 (65) 0.64 0.20 0.20 0.00
AP8 Methodology Availability 0 (9), 1 (91) 0.82 -0.09 -0.08 0.11
Average Approach 0.77 0.22 0.26 0.18
MD1 Methodology Completeness 0 (45), 1 (55) 0.69 0.38 0.46 0.24
MD2 Methodology Relevance 0 (26), 1 (74) 0.73 0.30 0.30 0.05
MD3 Methodology Target group 0 (25), 1 (75) 0.73 0.27 0.30 0.13
MD4 Methodology Existing Material 0 (35), 1 (65) 0.55 0.02 0.03 -0.13
MD5 Methodology Difficulties 0 (44), 1 (56) 0.56 0.12 0.19 0.25
MD6 Methodology Possibilities/restrictions 0 (56), 1 (44) 0.60 0.19 0.27 0.25
MD7 Methodology Details 0 (53), 1 (47) 0.49 -0.01 0.01 0.18
Average Methodology 0.62 0.18 0.22 0.14
SC1 Schedule Availability 0 (4), 1 (96) 0.96 0.49 0.57 0.04
SC2 Schedule Completeness 0 (4), 1 (96) 0.96 0.49 0.57 0.04
SC3 Schedule Block Description 0 (6), 1 (94) 0.91 0.24 0.28 -0.05
SC4 Schedule Realistic Relevance 0 (7), 1 (71), 2 (22) 0.90 0.21 0.36 0.25
Average Schedule 0.93 0.36 0.44 0.07
BI1 Bibliography Consistency 0 (19), 1 (81) 0.80 0.36 0.49 0.20
BI2 Key literature 0 (15), 1 (35), 2 (49) 0.76 0.03 0.35 0.64
Average Bibliography 0.78 0.19 0.42 0.42
AD1 Additional points 0 (97), 1 (2), 2 (1) 0.99 0.67 0.70 0.00
AD2 Negative points -2 (2), -1 (1), 0 (97) 0.96 -0.02 -0.03 -0.02
Average Additional points 0.98 0.32 0.34 -0.01
Average all criteria 0.79 0.20 0.27 0.16
Table 19: Quadratic weighted agreement (QWA) between Group 1 and Group 2 raters for each exposé criterion, with rubric-wise and overall averages. Columns show agreement (Agr), reliability ($\alpha$), correlation ($r$), and mean bias ($\Delta\mu$). Rows marked with have a highly skewed class distribution ($\max\_cp(c)\geq 90\%$).
Code Criterion Distribution (%) QWA $\alpha$ $r$ $\Delta\mu$
SC1 Summary available 0 (14), 1 (30), 2 (56) 0.90 0.49 0.61 0.11
SC2 Clarity 0 (10), 1 (38), 2 (53) 0.82 0.04 0.15 0.38
Average Structure and Clarity 0.86 0.27 0.38 0.25
LG1 Language 0 (6), 1 (94) 0.88 -0.06 -0.06 -0.07
LG2 Tone 0 (1), 1 (7), 2 (92) 0.98 0.23 0.39 -0.06
Average Language 0.93 0.08 0.17 -0.06
FU1 Learning Goals 0 (29), 1 (28), 2 (42) 0.83 0.51 0.50 -0.08
FU2 Success Criteria 0 (32), 1 (31), 2 (37) 0.79 0.35 0.36 0.05
Average Feed Up 0.81 0.43 0.43 -0.01
FB1 Knowledge of result 0 (5), 1 (95) 0.95 0.26 0.40 -0.05
FB2 Knowledge of correct response 0 (8), 1 (92) 0.90 -0.05 -0.05 -0.05
FB3 Knowledge about task constraints 0 (33), 1 (67) 0.65 0.07 0.08 0.07
FB4 Knowledge about concepts 0 (13), 1 (87) 0.82 -0.00 0.05 -0.12
FB5 Knowledge about mistakes 0 (8), 1 (92) 0.90 -0.05 -0.05 0.01
FB6 Self-feedback 0 (3), 1 (97) 0.97 -0.01 0.03
Average Feed Back 0.86 0.04 0.08 -0.02
FF1 Self-regulation 0 (6), 1 (94) 0.95 -0.02 -0.02 -0.03
FF2 Learning Skills 0 (39), 1 (61) 0.60 0.06 0.10 0.17
Average Feed Forward 0.78 0.02 0.04 0.07
CQ1 Correctness 0 (1), 1 (28), 2 (71) 0.91 0.14 0.16 0.12
CQ2 Actionability 0 (1), 1 (32), 2 (66) 0.92 0.22 0.21 0.08
CQ3 Argumentation 0 (10), 1 (90), 2 (0) 0.98 0.30 0.36 -0.06
Average Content Quality 0.94 0.22 0.24 0.05
ER1 Neglect -1 (16), 0 (84) 0.86 0.40 0.41 -0.05
ER2 Vague Critic -1 (7), 0 (93) 0.91 0.12 0.12 0.00
ER3 Out-of-Scope -1 (4), 0 (96) 0.93 -0.03 -0.03 -0.01
ER4 Missing Reference -1 (2), 0 (98) 0.96 -0.01 -0.04
ER5 Contradiction -1 (2), 0 (98) 0.98 -0.00 -0.01 0.00
Average Errors 0.93 0.10 0.12 -0.02
AD1 Additional points 0 (100), 1 (0) 1.00 0.00
AD2 Negative points -1 (2), 0 (98) 0.96 -0.01 -0.04
Average Additional points 0.98 -0.01 -0.02
Average all criteria 0.89 0.13 0.18 0.02
Table 20: Quadratic weighted agreement (QWA) between review Group 1 and Group 2 raters for each review criterion, with rubric-wise and overall averages. Columns show agreement (Agr), reliability ($\alpha$), correlation ($r$), and mean bias ($\Delta\mu$). Rows marked with have a highly skewed class distribution ($\max\_cp(c)\geq 90\%$).
Code Criterion Distribution (%) QWA $\alpha$ $r$ $\Delta\mu$
SC1 Summary available 0 (13), 1 (43), 2 (43) 0.91 0.60 0.62 0.00
SC2 Clarity 0 (12), 1 (48), 2 (40) 0.88 0.47 0.45 0.17
Average Structure and Clarity 0.89 0.53 0.54 0.08
LG1 Language 0 (4), 1 (96) 0.95 0.38 0.38 0.02
LG2 Tone 0 (1), 1 (12), 2 (88) 0.93 0.02 0.00 0.07
Average Language 0.94 0.20 0.19 0.04
FU1 Learning Goals 0 (28), 1 (28), 2 (44) 0.75 0.31 0.33 0.25
FU2 Success Criteria 0 (34), 1 (35), 2 (31) 0.72 0.13 0.15 0.23
Average Feed Up 0.74 0.22 0.24 0.24
FB1 Knowledge of result 0 (6), 1 (94) 0.88 -0.05 -0.06 -0.02
FB2 Knowledge of correct response 0 (11), 1 (89) 0.85 0.23 0.22 -0.02
FB3 Knowledge about task constraints 0 (37), 1 (63) 0.53 0.00 0.03 0.17
FB4 Knowledge about concepts 0 (13), 1 (87) 0.77 -0.00 -0.01 0.03
FB5 Knowledge about mistakes 0 (10), 1 (90) 0.87 0.27 0.27 0.03
FB6 Self-feedback 0 (4), 1 (96) 0.92 -0.03 -0.04 0.02
Average Feed Back 0.80 0.07 0.07 0.04
FF1 Self-regulation 0 (9), 1 (91) 0.82 -0.09 -0.10 -0.02
FF2 Learning Skills 0 (48), 1 (52) 0.57 0.14 0.14 0.10
Average Feed Forward 0.69 0.02 0.02 0.04
CQ1 Correctness 0 (0), 1 (39), 2 (61) 0.91 0.27 0.27 0.02
CQ2 Actionability 0 (2), 1 (34), 2 (63) 0.89 0.23 0.22 0.05
CQ3 Argumentation 0 (12), 1 (88), 2 (0) 0.94 -0.12 -0.13 0.00
Average Content Quality 0.91 0.13 0.12 0.02
ER1 Neglect -1 (15), 0 (85) 0.80 0.22 0.22 -0.03
ER2 Vague Critic -1 (5), 0 (95) 0.90 -0.04 -0.05 0.00
ER3 Out-of-Scope -1 (2), 0 (98) 0.95 -0.02 -0.02 -0.02
ER4 Missing Reference -1 (1), 0 (99) 0.98 0.00 0.02
ER5 Contradiction -1 (2), 0 (98) 0.95 -0.02 -0.02 -0.02
Average Errors 0.92 0.03 0.03 -0.01
AD1 Additional points 0 (100), 1 (0) 1.00 0.00
AD2 Negative points -1 (1), 0 (99) 0.98 0.00 0.02
Average Additional points 0.99 0.00 0.01
Average all criteria 0.86 0.13 0.14 0.04
Table 21: Quadratic weighted agreement (QWA) within Group 2 raters for each review criterion, with rubric-wise and overall averages. Columns show agreement (Agr), reliability ($\alpha$), correlation ($r$), and mean bias ($\Delta\mu$). Rows marked with have a highly skewed class distribution ($\max\_cp(c)\geq 90\%$).
G.2 Human–human agreement
Agreement reporting
We report raw quadratic weighted agreement (QWA), Krippendorff’s $\alpha$, Pearson correlation coefficient $r$, and simple mean difference (“mean bias”) between groups. QWA is the observed agreement between two raters on an ordinal scale, where disagreements are weighted by squared distance so that larger score differences are penalized more strongly than adjacent-category mismatches. Let $x_{i},y_{i}$ be two raters’ scores for item $i$ on an ordinal scale with discrete domain $\mathcal{D}=\{d_{\min},\dots,d_{\max}\}$ and span $\Delta=d_{\max}-d_{\min}>0$. We use the quadratic weight
$w(x_{i},y_{i})=\max\!\left(0,\,1-\left(\frac{x_{i}-y_{i}}{\Delta}\right)^{2}\right),$
and define raw quadratic weighted agreement as
$\mathrm{QWA}=\frac{1}{N}\sum_{i=1}^{N}w(x_{i},y_{i}),$
where $N$ is the number of co-rated items.
We use quadratic weighting because it preserves the magnitude of disagreement and remains interpretable under strong prevalence skew, a setting in which chance-corrected coefficients (e.g., $\kappa$ or $\alpha$) can be misleading. Accordingly, for highly skewed criteria (e.g., Number of pages, Template used, Negative points), we emphasize QWA.
$\alpha$ values range between -1 and 1, with 1 being perfect agreement, 0 being agreement by chance, and negative values indicating systematic disagreement; $r$ measures the correlation between a system’s predicted scores and the annotator-assigned scores; it ranges from -1 to 1. A positive (negative) value implies that the two sets of predictions are positively (negatively) correlated. Inter-annotator agreement results for exposé criteria are reported in Table 19, while agreement for review criteria is summarized in Tables 20 and 21.
G.3 Failure modes of chance-corrected agreement
This section interprets the agreement patterns observed in the exposé and review analyses (see Section 5.2 in the main paper). In particular, we focus on cases where chance-corrected reliability or correlation is negative, undefined, or unexpectedly low, despite high raw agreement, and relate these outcomes to prevalence effects, calibration differences between rater groups, and ambiguities in rubrics or training.
Prevalence effects under extreme class imbalance (“rare-event” criteria)
For highly skewed binary (or near-binary) criteria, the expected agreement under a chance model is already very high, so a handful of rare, non-overlapping deviations can yield $\alpha<0$ even when raw agreement is high. This explains the slightly negative values for Template used (ST1) and Negative points (AD2), where almost all submissions receive the default label, but the few exceptional ratings occur on different submissions. In such settings, Pearson’s $r$ is typically close to zero because the rating vectors are nearly constant, and mean bias $\Delta\mu$ helps diagnose whether the deviations are symmetric (ST1: $\Delta\mu\approx 0$) or slightly one-sided (AD2: small negative $\Delta\mu$ consistent with marginally more penalizing scores in one group).
Annotation guideline/interpretation issues (training or rubric ambiguity)
Some negative/undefined reliability values are attributable to concrete, identifiable rubric failures. For Metadata available (ME2), metadata was present for all submissions; all disagreements originated from a single rater in Group 2, indicating a training/interpretation error (i.e., the rater did not apply the intended check). This also leads to an undefined Pearson correlation because one group’s ratings are effectively constant (zero variance). For Template used (ST1), the few “0” assignments were triggered by superficial deviations (e.g., renamed chapter headings or bibliography compilation issues) despite the underlying template being used, revealing an unclear rubric description about what counts as a template violation. For Methodology Availability (AP8), we found edge cases where the methodology is present only as a very short list; these items were inconsistently judged as available vs. unavailable across groups, suggesting that the criterion definition is underspecified (i.e., availability vs. sufficiency/quality).
Systematic calibration shifts and low signal in subjective criteria
For Anker (MO2), the negative (near-zero) $\alpha$ is consistent with a systematic calibration shift between groups: Group 2 assigns lower scores on average (positive mean bias $\Delta\mu=0.309$), and disagreements are predominantly adjacent-category mismatches (e.g., $2$ vs. $1$), which remain relatively benign under quadratic weighting but still depress chance-corrected reliability. For Problem relevance (MO4), negative coefficients are driven by sparse and subjective “0” decisions: the negative label is rare and tends to be applied to different submissions across groups.
Prevalence effects under extreme class imbalance in review scoring
A closely related prevalence-driven failure mode appears in review scoring for error and penalty criteria (Tables 20 and 21). These criteria are effectively rare-event detection tasks on a highly skewed label space: the non-penalizing default (typically 0) dominates, while penalizing assignments (e.g., 1-1 in Errors (ER)) occur only rarely. As a result, chance-corrected reliability and Pearson correlation become unstable or uninformative. A small number of non-overlapping penalizations can yield $\alpha\leq 0$ despite high raw agreement, and near-constant rating vectors drive $r$ toward zero or render it undefined when one rater shows zero variance.
Overall, we emphasize quadratic-weighted raw agreement as our primary agreement signal and use $\alpha$, $r$, $\Delta\mu$, and the pooled distributions to (i) flag prevalence-induced artifacts, (ii) diagnose calibration differences, and (iii) identify rubric/training issues that can be addressed in the future.
Qwen3 (80B) Llama 3.3 (70B) GPT OSS (120B) GPT OSS (20B)
Code $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$
LA1 0.83 0.85 0.85 0.91 0.78 0.84 0.77 0.83 0.68 0.59 0.73 0.67 0.72 0.57 0.77 0.64
LA2 0.87 0.88 0.76 0.85 0.87 0.87 0.75 0.75 0.88 0.90 0.89 0.86 0.82 0.87 0.91 0.81
Language 0.85 0.87 0.80 0.88 0.82 0.85 0.76 0.79 0.78 0.75 0.81 0.76 0.77 0.72 0.84 0.73
ME1 0.91 0.92 0.82 0.86 0.91 0.90 0.84 0.87 0.91 0.84 0.87 0.84 0.82 0.81 0.90 0.86
Metadata 0.91 0.92 0.82 0.86 0.91 0.90 0.84 0.87 0.91 0.84 0.87 0.84 0.82 0.81 0.90 0.86
MO1 0.87 0.87 0.82 0.82 0.87 0.84 0.82 0.82 0.84 0.73 0.78 0.60 0.84 0.71 0.78 0.69
MO2 0.86 0.88 0.74 0.76 0.86 0.86 0.74 0.74 0.92 0.87 0.82 0.76 0.90 0.86 0.85 0.74
MO3 0.96 0.98 0.91 0.93 0.98 0.98 0.93 0.93 0.98 0.98 0.93 0.93 0.98 0.98 0.93 0.93
MO4 0.38 0.82 0.42 0.84 0.75 0.82 0.76 0.84 0.82 0.82 0.84 0.84 0.82 0.82 0.84 0.84
MO5 0.75 0.75 0.56 0.56 0.75 0.75 0.56 0.56 0.76 0.75 0.55 0.56 0.73 0.75 0.55 0.56
MO6 0.87 0.85 0.79 0.78 0.86 0.87 0.78 0.79 0.92 0.89 0.85 0.85 0.90 0.87 0.81 0.81
MO7 0.67 0.69 0.40 0.42 0.67 0.67 0.40 0.40 0.75 0.65 0.47 0.71 0.67 0.76 0.40 0.49
Motivation 0.77 0.83 0.66 0.73 0.82 0.83 0.71 0.72 0.85 0.81 0.75 0.75 0.83 0.82 0.74 0.72
AP1 0.89 0.91 0.87 0.84 0.91 0.91 0.84 0.84 0.93 0.85 0.85 0.82 0.93 0.89 0.89 0.85
AP2 0.86 0.85 0.76 0.71 0.87 0.84 0.79 0.70 0.89 0.88 0.81 0.77 0.86 0.86 0.91 0.74
AP3 0.75 0.70 0.64 0.60 0.75 0.72 0.72 0.65 0.78 0.77 0.73 0.66 0.85 0.75 0.87 0.67
AP4 0.15 0.78 0.11 0.62 0.80 0.78 0.64 0.62 0.82 0.76 0.65 0.67 0.84 0.82 0.67 0.65
AP5 0.62 0.76 0.49 0.56 0.80 0.76 0.60 0.56 0.84 0.75 0.56 0.62 0.82 0.84 0.62 0.67
AP6 0.83 0.77 0.76 0.66 0.80 0.76 0.74 0.63 0.81 0.84 0.88 0.80 0.78 0.83 0.88 0.82
AP7 0.16 0.65 0.15 0.65 0.64 0.65 0.64 0.65 0.67 0.71 0.71 0.60 0.56 0.75 0.65 0.71
AP8 0.96 0.96 0.85 0.85 0.96 0.96 0.85 0.85 0.96 0.96 0.85 0.85 0.96 0.96 0.85 0.85
Approach 0.65 0.80 0.58 0.69 0.82 0.80 0.73 0.69 0.84 0.82 0.76 0.72 0.82 0.84 0.79 0.75
MD1 0.67 0.67 0.44 0.44 0.67 0.67 0.44 0.44 0.69 0.64 0.45 0.58 0.65 0.78 0.49 0.58
MD2 0.73 0.76 0.69 0.71 0.75 0.76 0.69 0.71 0.76 0.75 0.71 0.69 0.76 0.76 0.71 0.71
MD3 0.80 0.82 0.67 0.69 0.80 0.82 0.67 0.69 0.85 0.78 0.73 0.69 0.87 0.78 0.75 0.73
MD4 0.56 0.58 0.73 0.71 0.56 0.58 0.73 0.71 0.62 0.64 0.71 0.73 0.60 0.58 0.73 0.67
MD5 0.71 0.69 0.45 0.44 0.71 0.71 0.49 0.45 0.78 0.64 0.56 0.64 0.82 0.70 0.71 0.70
MD6 0.58 0.56 0.33 0.31 0.58 0.56 0.36 0.31 0.67 0.73 0.45 0.55 0.64 0.65 0.56 0.55
MD7 0.56 0.56 0.38 0.38 0.56 0.56 0.38 0.38 0.58 0.60 0.44 0.45 0.60 0.60 0.42 0.42
Methodology 0.66 0.66 0.53 0.52 0.66 0.67 0.54 0.53 0.71 0.68 0.58 0.62 0.71 0.70 0.62 0.62
SC1 0.98 0.98 0.95 0.95 0.98 0.98 0.95 0.95 0.96 0.96 0.96 0.96 0.98 0.96 0.95 0.96
SC2 0.98 0.98 0.95 0.95 0.98 0.98 0.95 0.95 0.98 0.95 0.95 0.91 0.98 0.98 0.95 0.95
SC3 0.91 0.91 0.96 0.96 0.91 0.91 0.96 0.96 0.91 0.93 0.96 0.98 0.91 0.91 0.96 0.96
SC4 0.85 0.82 0.85 0.78 0.90 0.78 0.95 0.72 0.85 0.83 0.88 0.81 0.85 0.77 0.88 0.74
Schedule 0.93 0.92 0.93 0.91 0.94 0.91 0.95 0.89 0.93 0.92 0.94 0.92 0.93 0.91 0.93 0.90
BI1 0.93 0.72 0.72 0.59 0.91 0.91 0.71 0.71 0.27 0.13 0.40 0.33 0.20 0.49 0.40 0.55
BI2 0.90 0.92 0.68 0.73 0.90 0.88 0.71 0.66 0.86 0.91 0.80 0.72 0.91 0.88 0.75 0.70
Bibliography 0.91 0.82 0.70 0.66 0.91 0.90 0.71 0.68 0.57 0.52 0.60 0.52 0.56 0.69 0.57 0.62
All criteria 0.75 0.80 0.66 0.70 0.81 0.80 0.71 0.70 0.81 0.77 0.74 0.72 0.79 0.79 0.75 0.73
Table 22: Quadratic-weighted agreement between LLM scores and human ratings for exposé criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Qwen3 (80B) Llama 3.3 (70B) GPT OSS (120B) GPT OSS (20B)
Code $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$ $A^{\text{all}}_{\text{G1}}$ $A^{\text{ind}}_{\text{G1}}$ $A^{\text{all}}_{\text{G2}}$ $A^{\text{ind}}_{\text{G2}}$
SC1 0.83 0.83 0.85 0.84 0.84 0.85 0.88 0.89 0.85 0.83 0.87 0.87 0.86 0.88 0.88 0.91
SC2 0.92 0.92 0.80 0.81 0.90 0.86 0.81 0.80 0.91 0.86 0.83 0.85 0.84 0.85 0.85 0.83
Structure and Clarity 0.88 0.87 0.82 0.83 0.87 0.85 0.84 0.85 0.88 0.84 0.85 0.86 0.85 0.87 0.87 0.87
LG1 0.90 0.91 0.98 0.97 0.91 0.88 0.97 0.94 0.86 0.67 0.85 0.62 0.75 0.69 0.75 0.64
LG2 0.98 0.98 0.99 0.98 0.98 0.98 0.99 0.99 0.98 0.99 0.99 1.00 0.96 0.97 0.98 1.00
Language 0.94 0.94 0.99 0.97 0.94 0.93 0.98 0.97 0.92 0.83 0.92 0.81 0.85 0.83 0.87 0.82
FU1 0.75 0.70 0.74 0.69 0.74 0.78 0.72 0.73 0.64 0.63 0.57 0.55 0.66 0.55 0.60 0.48
FU2 0.75 0.71 0.73 0.69 0.76 0.78 0.72 0.75 0.64 0.66 0.62 0.65 0.58 0.61 0.59 0.60
Feed Up 0.75 0.71 0.73 0.69 0.75 0.78 0.72 0.74 0.64 0.64 0.59 0.60 0.62 0.58 0.59 0.54
FB1 0.94 0.94 0.99 0.99 0.94 0.92 0.99 0.96 0.23 0.10 0.20 0.06 0.25 0.09 0.22 0.05
FB2 0.92 0.93 0.97 0.96 0.91 0.93 0.95 0.96 0.24 0.09 0.19 0.05 0.08 0.10 0.04 0.04
FB3 0.79 0.79 0.73 0.73 0.79 0.78 0.73 0.74 0.77 0.62 0.73 0.63 0.75 0.59 0.69 0.58
FB4 0.84 0.84 0.96 0.96 0.84 0.84 0.96 0.96 0.83 0.74 0.93 0.78 0.84 0.84 0.96 0.91
FB5 0.95 0.95 0.94 0.94 0.95 0.94 0.94 0.92 0.95 0.92 0.94 0.92 0.95 0.76 0.94 0.79
FB6 1.00 0.99 0.97 0.96 0.99 0.99 0.98 0.98 0.96 0.83 0.95 0.82 0.34 0.41 0.37 0.42
Feed Back 0.91 0.91 0.93 0.92 0.90 0.90 0.93 0.92 0.66 0.55 0.66 0.54 0.54 0.47 0.54 0.46
FF1 0.96 0.96 0.99 0.99 0.96 0.96 0.99 0.99 0.96 0.90 0.99 0.92 0.91 0.71 0.90 0.70
FF2 0.78 0.78 0.61 0.61 0.78 0.78 0.61 0.61 0.75 0.37 0.62 0.50 0.78 0.36 0.71 0.51
Feed Forward 0.87 0.87 0.80 0.80 0.87 0.87 0.80 0.80 0.86 0.63 0.81 0.71 0.84 0.53 0.80 0.60
CQ1 0.96 0.94 0.92 0.92 0.94 0.88 0.93 0.86 0.92 0.95 0.90 0.92 0.84 0.90 0.84 0.90
CQ2 0.94 0.94 0.92 0.92 0.93 0.94 0.93 0.91 0.94 0.94 0.92 0.92 0.92 0.93 0.93 0.92
CQ3 0.77 0.98 0.80 0.99 0.98 0.98 0.99 0.99 0.98 0.96 0.99 0.96 0.98 0.96 0.99 0.97
Content Quality 0.89 0.95 0.88 0.94 0.95 0.93 0.95 0.92 0.95 0.95 0.94 0.93 0.91 0.93 0.92 0.93
ER1 0.83 0.63 0.88 0.68 0.83 0.44 0.88 0.45 0.75 0.65 0.77 0.70 0.80 0.82 0.83 0.87
ER2 0.92 0.74 0.92 0.75 0.93 0.65 0.95 0.61 0.89 0.91 0.89 0.91 0.86 0.94 0.88 0.94
ER3 0.96 0.91 0.97 0.92 0.96 0.93 0.97 0.94 0.96 0.92 0.97 0.93 0.96 0.92 0.97 0.93
ER4 0.80 0.25 0.84 0.26 0.94 0.17 0.98 0.13 0.42 0.32 0.38 0.32 0.36 0.36 0.38 0.34
ER5 0.99 0.95 0.99 0.95 0.99 0.94 0.99 0.94 0.99 0.97 0.99 0.97 0.99 0.95 0.99 0.95
Errors 0.90 0.69 0.92 0.71 0.93 0.63 0.95 0.62 0.80 0.75 0.80 0.77 0.79 0.80 0.81 0.81
All criteria 0.89 0.84 0.89 0.84 0.90 0.83 0.90 0.82 0.79 0.72 0.78 0.72 0.74 0.69 0.74 0.69
Table 23: Quadratic-weighted agreement between LLM scores and human ratings for review criteria. Columns report $A_{G1}$ (Group 1 vs. LLM) and $A_{G2}$ (Group 2 vs. LLM) for the prompting variants shown (all, ind). Bold/underlined cells mark column-wise maxima/minima. Average rows summarize agreement within rubrics and across all criteria.
Appendix H Experimental Details
Models
We evaluate a diverse set of recent open-weight instruction-tuned LLMs that cover different architectures and scales. Our selection includes (i) a large dense decoder-only Transformer from the Llama 3.3 family (meta-llama/Llama-3.3-70B-Instruct; 70B parameters) Team (2024), (ii) a high-sparsity Mixture-of-Experts (MoE) model with long-context support (Qwen/Qwen3-Next-80B-A3B-Instruct; 80B total parameters with 3B active) Team (2025) and an independent MoE reasoning model in two flavors, (iii) the high reasoning model (openai/gpt-oss-120b, 117B parameters with 5.1B active parameters) and (iv) the smaller low latency variant (openai/gpt-oss-20b, 21B parameters with 3.6B active parameters) OpenAI (2025).
All models are fully open-weight and can be run on two local NVIDIA A100 80GB GPUs, which is important for classroom deployment where data protection and limited consent often require on-premise inference instead of cloud-based APIs. Inference is performed using the Hugging Face Transformers library [^15] https://huggingface.co/docs/transformers with deterministic greedy decoding (do_sample = False).
Prompts
For both tasks, we use a unified prompting strategy (Appendix I) with a system message that (i) assigns the model the role of an expert evaluator and (ii) enumerates the relevant criteria, including their names, short descriptions, allowed integer score ranges, and the verbal descriptions associated with each score level (Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric). The model is required to return a single JSON object with fields criterion_name, assigned_score, and a short justification. The additional justifications are requested to explain the decision-making process, as it is crucial in an environment where people are affected by the LLM outcome Kolářová and Schmude (2026).
Criteria excluded from the LLM experiments
The exposé and review rubrics contain structural or format-related checks (e.g., LaTeX template usage, page limits) and content-related criteria. We exclude the first one, as the LLM cannot successfully assess it because the generated PDF is needed for correct evaluation. We also exclude additional points that can be given by the human evaluator, as these occur only in isolated cases and are intended for exceptional circumstances. The following exposé criteria are excluded: Metadata available; Template used; Number of pages; Additional points; Negative points. The following review criteria are excluded: Additional points; Negative points.
Appendix I Prompt Templates
This appendix provides the exact prompt formats used in all baseline experiments. The rubric text included in the templates is automatically generated from the released JSON rubric files, ensuring that the prompts align with the annotation scheme used in Exposía.
For both tasks, we use a unified prompt structure. The system message (i) assigns the model the role of an expert evaluator and (ii) enumerates the relevant criteria, including their names, short descriptions, allowed integer score ranges, and the verbal descriptions associated with each score level (Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric). The model is required to return a JSON list with one object per criterion, containing the fields criterion_name, assigned_score, and a short justification. The additional justifications are requested to explain the decision-making process, as it is crucial in an environment where people are affected by the LLM outcome Kolářová and Schmude (2026).
I.1 Expose Scoring Prompt
The user message provides the topic, the full exposé text, and the bibliography.
System message You are an expert evaluator specializing in assessing student exposes. Your task is to carefully evaluate the given expose based on the specified grading criteria. The grading criteria are as follows: ## <Rubric Name> <Rubric description> ### Criterion: <Criterion Name> Description: <Criterion description> Allowed scores: - <Points> points: <Description> ... ## <Next Rubric> ... You must: 1. Evaluate the expose **criterion by criterion**. 2. For **every** listed criterion, assign exactly **one** score using **only** the scoring options and point values provided. 3. Justify each score with a clear, concise and **short** explanation. Your response MUST be a list of dictionaries in **valid JSON** schema: [ { "criterion_name": "<the EXACT criterion name from the criteria list, without any other words>", "assigned_score": "<an INTEGER, one of the allowed point values for this criterion>", "justification": "<SHORT justification of why you decided to give this score>" }, ... ] Constraints: - Return **one object per criterion**; do not skip any criteria. - Use only the criteria and scoring options that are explicitly provided. - Return a response that consists solely of valid JSON, with no additional explanatory text before or after it. - Make sure the JSON is syntactically valid (no trailing commas, correct quotation marks, etc.).
User message The topic is: "<TOPIC>". Here is the expose you have to grade: <EXPOSE TEXT> Bibliography: <BIBLIOGRAPHY>
I.2 Review Scoring Prompt
The user message provides the full written review text.
System message You are an expert evaluator specializing in assessing the quality of student peer reviews. Your task is to carefully evaluate the reviewer's performance based on the specified grading criteria. Grading criteria: ## <Rubric Name> <Rubric description> ### Criterion: <Criterion Name> Description: <Criterion description> Allowed scores: - <Points> points: <Description> ... ## <Next Rubric> ... You must: 1. Evaluate the peer review **criterion by criterion**. 2. For **every** listed criterion, assign exactly **one** score using **only** the scoring options and point values provided for that criterion. 3. Base your evaluation strictly on the quality of the full review text. 4. Justify each score with a clear, concise and **short** explanation. Output format: Return a list of objects following this **valid JSON** schema: [ { "criterion_name": "<EXACT criterion name from the criteria list>", "assigned_score": <INTEGER, one of the allowed point values for this criterion>, "justification": "<SHORT justification of why you decided to give this score>" }, ... ] Constraints: - Return **one object per criterion**; do not skip any criteria. - Use only the criteria and scoring options that are explicitly provided. - The response must consist solely of valid JSON, with no additional explanatory text. - Make sure the JSON is syntactically valid (no trailing commas, correct quotation marks, etc.).
User message Here are the materials of the peer review you have to grade. === Full written text of the review === <FULL_REVIEW_TEXT>
Appendix J Additional Experimental Results
Agreement by expertise level
Table 4 reports human–LLM agreement (QWA) for criterion-level scoring, grouped by criterion expertise. Across both tasks, agreement decreases monotonically with expertise, mirroring the human–human reference. For exposé scoring, low-expertise criteria (Exp. 0; three criteria) yield high agreement for the strongest models under combined prompting (e.g., Llama 3.3: $\overline{A}_{\text{all}}{=}0.90$; Qwen3: $0.91$), whereas Exp. 2 criteria are substantially harder (Llama 3.3: $0.63$; Qwen3: $0.57$); humans show the same drop in the Human column (0.89 at Exp. 0 vs. 0.64 at Exp. 2), consistent with higher-expertise criteria being more subjective. Aggregated over all 31 exposé criteria, model agreement is close to the human baseline (Human: 0.76), and combined prompting is typically comparable or slightly better than single-criterion prompting, with Qwen3 as the main exception ($\overline{A}_{\text{ind}}{=}0.75$ vs. $\overline{A}_{\text{all}}{=}0.72$). The comparatively lower Exp. 0 values for GPT-OSS in the aggregate are explained by the small size of this bin and a single outlier criterion: Table 22 shows that GPT-OSS matches humans very well on the schedule checks (SC1/SC3; QWA $\approx 0.91$$0.98$), but performs unusually poorly on BI1 (Bibliography Consistency; e.g., GPT-OSS 120B: $A^{\text{all}}_{G1}{=}0.27$, $A^{\text{all}}_{G2}{=}0.40$), which dominates the Exp. 0 average. For review scoring, overall agreement is higher (Human: 0.88 over 22 criteria), and Llama 3.3 and Qwen3 reach near-human alignment under combined prompting (Llama 3.3: $\overline{A}_{\text{all}}{=}0.90$; Qwen3: $0.89$); the expertise effect persists, with Exp. 2 criteria yielding lower agreement than Exp. 1.
Figure 7: Human–LLM agreement (QWA) for review scoring by expertise level. The dotted line shows human-human QWA.
Figure 8: Group-asymmetry in human–LLM agreement by review rubric. Each point shows the difference in agreement between an LLM and human raters (Group 1 (G1) vs. Group 2 (G2)). The vertical line at $\Delta=0$ indicates equal agreement of the LLM with human G1 and G2 raters. Points to the right indicate the model agrees more with G1 raters than with G2 raters.
Figure 9: Runtime per submission for criterion-based assessments across models. Bars show mean seconds per submission with standard deviation bars. Within each task, All denotes combined prompting (light fill) and Ind denotes single-criterion prompting (dark fill). Review runs are hatched; Exposé runs are unhatched. Ratios show $\times(\mathrm{Ind}/\mathrm{All})$, highlighting the additional latency introduced by single-criterion prompting.
Scalability and classroom deployment
Across all models and both tasks, combined prompting is more efficient in most cases than single-criterion prompting: it reduces mean runtime per submission by about $1.3\times$$5.3\times$ and avoids the linear cost growth from issuing one model call per criterion (Figure 9). Single-criterion prompting also substantially increases total token processing because the full input is repeated across prompts, with particularly large overhead for long exposés and the review setting. These efficiency differences become especially consequential under the observed course usage dynamics. As shown by the behavioural interaction logs (Appendix Figure 6), student activity is strongly deadline-driven, with a pronounced spike in actions immediately before the review deadline, whereas instructor activity is relatively steady. This implies that in practice, LLM-based scoring must be provisioned for short periods of peak demand rather than average-day usage. As such, the overhead of single-criterion prompting is particularly problematic, while combined prompting better supports batching and throughput during deadline-induced bursts. Consequently, combined prompting is the recommended operating point for classroom deployment under constrained on-premise resources, with model choice primarily determining the remaining latency–quality trade-off. Runtime and token usage are reported in Appendix K (Tables 2425).
Operational latency under peak-demand grading windows
In classroom deployment under constrained on-premise resources, using LLM scores as instructor support (e.g., batch pre-scoring or rubric-aligned decision support) introduces a practical scheduling choice: scores can either be pre-computed ahead of grading, or compute capacity must be provisioned to serve requests on demand throughout the grading period. Under deadline-driven bursts (Appendix Figure 6), it can translate directly into waiting times, because even combined prompting incurs minute-level latency per submission and shows substantial min–max spread (Tables 2425). To illustrate, consider a cohort of N=600N{=}600 exposés scored on the same two-GPU setup used in our experiments. For exposés, the fastest combined configuration in Table 24 (GPT OSS 20B, All) yields $600\times 110.2\text{s}\approx 18.4$ hours on average; larger models further increase this budget (e.g., Qwen3 80B, All: $\approx 71.7$ hours). For student reviews, the volume is larger ($2N$ submissions); in the worst-case combined configuration, runtimes of $\approx 289.6$s per review (Table 25) translate into $\approx 96.5$ hours of wall-clock time, making peak-time live scoring particularly susceptible to queueing delays.
Finally, beyond latency, we observed occasional out-of-memory failures for unusually long inputs (e.g., exposés with many and very long comments), reinforcing the need for robust deployment to plan around worst-case input lengths (e.g., advance computation, batching, and input-length safeguards) rather than relying on unconstrained live scoring near deadlines.
Is review text enough for review assessment?
Overall, we did not find an indication that review scoring requires context beyond the review text. In supplementary experiments, we augmented the review input jointly with additional context from the exposé and inline comments, but did not observe consistent improvements over the review-only setting.
Appendix K Runtime and Token Consumption
In this appendix, we focus exclusively on inference-time cost (runtime and token usage) under these fixed conditions. Prompting variants (combined vs. single-criterion) are defined in Section 6.1. For each model/condition, we measure wall-clock runtime both per full submission (s/assessment) and per model call (s/prompt). Token counts are computed with the respective model tokenizer on the fully formatted prompt (system + user messages) and the generated JSON response, and reported as tokens/prompt split into tokens/input and tokens/output. All values are aggregated across evaluation instances and reported as mean (SD) and [min–max], capturing variation in input length.
LLM scoring throughput: combined prompting is essential
Tables 24 and 25 report runtime and token usage for exposé and review scoring, respectively. Across both tasks, combined prompting is essential for throughput: it requires a single forward pass per submission and therefore yields the lowest latency (110–430s on average across models for exposé scoring). In contrast, single-criterion prompting multiplies calls per submission, increasing total runtime to 284–1133s on average for exposés, with worst cases up to 23 minutes. The same multiplicative effect appears in token processing: although single-criterion outputs are shorter per call, repeating the full input across criteria yields an order-of-magnitude increase in total processed tokens per submission (approximately $13\times$ in our setup). Review scoring exhibits the same trend (Table 25), with combined prompting consistently yielding lower latency and substantially fewer processed tokens per submission than single-criterion prompting.
Model Mode s/assessment s/prompt tokens/prompt tokens/input tokens/output
GPT OSS (20B) All 110.20 ($\pm$27.69) 110.20 ($\pm$27.69) 8,010.67 ($\pm$1,018.78) 5,587.85 ($\pm$932.59) 2,422.82 ($\pm$629.58)
[min–max] [84.88–289.55] [84.88–289.55] [6,384.00–11,178.00] [4,092.00–8,765.00] [1,874.00–6,489.00]
Llama 3.3 (70B) All 182.10 ($\pm$12.79) 182.10 ($\pm$12.79) 7,181.78 ($\pm$973.95) 5,606.69 ($\pm$947.16) 1,575.09 ($\pm$105.46)
[min–max] [163.18–212.11] [163.18–212.11] [5,594.00–10,311.00] [4,091.00–8,839.00] [1,399.00–1,858.00]
GPT OSS (120B) All 258.01 ($\pm$13.74) 258.01 ($\pm$13.74) 8,227.11 ($\pm$915.79) 5,679.85 ($\pm$932.59) 2,547.25 ($\pm$148.73)
[min–max] [227.52–304.39] [227.52–304.39] [6,692.00–11,292.00] [4,184.00–8,857.00] [2,265.00–3,067.00]
Llama 3.3 (70B) Ind 284.30 ($\pm$26.54) 9.17 ($\pm$1.37) 3,228.31 ($\pm$948.58) 3,162.76 ($\pm$947.61) 65.55 ($\pm$9.79)
[min–max] [237.56–370.23] [5.43–14.43] [1,656.00–6,555.00] [1,613.00–6,490.00] [42.00–117.00]
Qwen3 (80B) All 430.37 ($\pm$15.13) 430.37 ($\pm$15.13) 7,473.75 ($\pm$1,046.99) 5,719.84 ($\pm$1,024.62) 1,753.91 ($\pm$64.05)
[min–max] [394.29–460.62] [394.29–460.62] [5,775.00–11,079.00] [4,126.00–9,301.00] [1,602.00–1,874.00]
GPT OSS (20B) Ind 582.36 ($\pm$103.30) 18.79 ($\pm$20.85) 3,553.87 ($\pm$1,047.56) 3,162.31 ($\pm$933.03) 391.56 ($\pm$478.39)
[min–max] [452.01–878.16] [7.34–357.11] [1,821.00–11,982.00] [1,633.00–6,431.00] [123.00–8,091.00]
Qwen3 (80B) Ind 862.02 ($\pm$36.97) 27.81 ($\pm$3.81) 3,359.56 ($\pm$1,027.13) 3,275.90 ($\pm$1,025.04) 83.66 ($\pm$15.20)
[min–max] [789.80–946.85] [19.10–41.08] [1,705.00–7,050.00] [1,648.00–6,952.00] [51.00–138.00]
GPT OSS (120B) Ind 1133.41 ($\pm$89.76) 34.35 ($\pm$6.90) 3,383.58 ($\pm$932.15) 3,160.07 ($\pm$933.04) 223.51 ($\pm$67.47)
[min–max] [869.61–1398.68] [19.20–97.15] [1,771.00–6,902.00] [1,627.00–6,431.00] [100.00–854.00]
Table 24: Runtime and token usage for LLM scoring. Entries report mean ($\pm$ SD) and a second row with [min–max]. Mode: All = combined prompting; Ind = single-criterion prompting. Token statistics are reported per prompt (total, input, output).
Model Mode s/assessment s/prompt tokens/prompt tokens/input tokens/output
GPT OSS (20B) All 80.25 ($\pm$20.05) 80.25 ($\pm$20.05) 4,465.78 ($\pm$606.62) 2,539.31 ($\pm$402.22) 1,926.47 ($\pm$472.59)
[min–max] [60.56–345.62] [60.56–345.62] [3,641.00–10,957.00] [2,117.00–6,330.00] [1,449.00–8,091.00]
Llama 3.3 (70B) All 113.76 ($\pm$8.10) 113.76 ($\pm$8.10) 3,574.05 ($\pm$434.77) 2,513.88 ($\pm$404.74) 1,060.17 ($\pm$73.19)
[min–max] [98.39–135.10] [98.39–135.10] [3,041.00–7,464.00] [2,088.00–6,329.00] [920.00–1,264.00]
Llama 3.3 (70B) Ind 142.31 ($\pm$11.27) 6.47 ($\pm$0.95) 982.85 ($\pm$406.97) 924.60 ($\pm$405.24) 58.25 ($\pm$8.26)
[min–max] [121.03–210.90] [4.26–11.25] [518.00–4,845.00] [471.00–4,774.00] [40.00–93.00]
GPT OSS (120B) All 194.64 ($\pm$14.13) 194.64 ($\pm$14.13) 4,485.63 ($\pm$447.10) 2,540.12 ($\pm$403.33) 1,945.51 ($\pm$146.74)
[min–max] [167.25–252.35] [167.25–252.35] [3,823.00–8,242.00] [2,117.00–6,330.00] [1,670.00–2,555.00]
Qwen3 (80B) All 289.58 ($\pm$13.70) 289.58 ($\pm$13.70) 3,705.31 ($\pm$483.11) 2,567.16 ($\pm$460.99) 1,138.15 ($\pm$50.89)
[min–max] [254.40–359.69] [254.40–359.69] [3,087.00–7,550.00] [2,066.00–6,295.00] [1,004.00–1,361.00]
GPT OSS (20B) Ind 343.64 ($\pm$81.03) 15.62 ($\pm$15.37) 1,316.56 ($\pm$561.36) 955.08 ($\pm$402.72) 361.47 ($\pm$363.39)
[min–max] [240.61–792.63] [4.91–344.67] [687.00–9,596.00] [505.00–4,779.00] [108.00–8,091.00]
Qwen3 (80B) Ind 485.12 ($\pm$25.59) 22.05 ($\pm$2.33) 970.92 ($\pm$407.79) 902.12 ($\pm$405.52) 68.80 ($\pm$9.00)
[min–max] [421.38–576.16] [15.73–40.33] [516.00–4,822.00] [449.00–4,740.00] [45.00–145.00]
GPT OSS (120B) Ind 512.82 ($\pm$88.18) 23.31 ($\pm$9.87) 1,171.90 ($\pm$420.31) 955.08 ($\pm$402.72) 216.81 ($\pm$96.63)
[min–max] [375.80–1048.64] [8.89–465.17] [634.00–5,883.00] [505.00–4,779.00] [85.00–4,827.00]
Table 25: Runtime and token usage for LLM scoring. Entries report mean ($\pm$ SD) and a second row with [min–max]. Mode: All = combined prompting; Ind = single-criterion prompting. Token statistics are reported per prompt (total, input, output).
Appendix L Behaviour Data: Interaction Logs and Semester Dynamics
To complement the textual artefacts in Exposía, we also record fine-grained behavioural interaction logs from the review platform used in the course.
Temporal dynamics across the semester
Figure 6 shows daily action counts for students and instructors. Two qualitative patterns stand out.
First, student activity is strongly deadline-driven: actions rise sharply in the days immediately preceding the review deadline (red dashed line). In contrast, instructor activity is comparatively smoother and distributed across the semester, consistent with instructional work that is scheduled and paced over time rather than concentrated in the final days.
Second, activity drops to near-zero during the winter break (hatched region), including the Christmas/New Year period, and then resumes afterward.
Implications for LLM-assisted assessment workflows
The pronounced pre-deadline student spike (Figure 6) suggests that any LLM-based support integrated into the reviewing/assessment pipeline (e.g., rubric guidance, feedback scaffolding, or automated checks) will face highly non-uniform demand. Compute budgets and latency constraints should be evaluated under peak load conditions rather than on average-day usage.
Limitations
These behavioural patterns require particular caution because they reflect only partial observation. Not all students consented to behavioural logging, so the curves describe only the consenting subset and may not generalize to the full cohort. The instructor group is also small (n=9n{=}9), limiting statistical power and making aggregates sensitive to individual work styles. Moreover, instructor activity is not directly comparable to student activity because instructors carried a substantially higher workload: while students completed only two reviews without grading, instructors collectively reviewed across the cohort and additionally assigned criterion-based grades. Finally, logs capture only on-platform actions; substantial reviewing work may occur offline (e.g., reading drafts externally or drafting feedback elsewhere). Thus, peaks and troughs should be interpreted as platform activity rather than complete measures of effort or time-on-task.